### v1

# Lie symmetry analysis of the two-Higgs-doublet model field equations

Lie symmetry analysis of the two-Higgs-doublet model field equations
M. Aa. Solberg [^1] E-mail: marius.solberg@ntnu.no
Department of Structural Engineering,
Norwegian University of Science and Technology,
Trondheim, Norway
We apply Lie symmetry analysis of partial differential equations (PDEs) to the Euler-Lagrange equations of the two-Higgs-doublet model (2HDM), to determine its scalar Lie point symmetries. A Lie point symmetry is a structure-preserving transformation of the spacetime variables and the fields of the model, which is also continuous and connected to the identity. Symmetries of PDEs may in general be divided into strict variational symmetries, divergence symmetries and non-variational symmetries, where the first two are collectively referred to as variational symmetries. Variational symmetries are usually preserved under quantization, and variational Lie symmetries yield conservation laws. We demonstrate that there are no scalar Lie point divergence symmetries or non-variational Lie point symmetries in the 2HDM, and re-derive its well-known strict variational Lie point symmetries, thus confirming the consistency of our implementation of Lie’s method. Moreover, we prove three general results which may simplify Lie symmetry calculations for a wide class of particle physics models. Lie symmetry analysis of PDEs is a broadly applicable method for determining Lie symmetries. As demonstrated here by example, it can be applied to models with many variables, parameters and reparametrization freedom, while any missing discrete symmetries may be identified through the automorphism groups of the resulting Lie symmetry algebras.
1 Introduction
Theories with extended scalar sectors beyond the Standard Model (SM) Higgs boson, discovered at CERN in 2012 [1, 2], may explain the matter-antimatter asymmetry of the Universe [3, 4, 5] and accommodate dark matter candidates [6, 7, 8, 9]. New scalars fields can be introduced in several ways, but the most phenomenologically viable and studied implementations are additional complex, scalar $\mathsf{SU}(2)_{L}$ doublets or scalar gauge singlets [10, 11, 12]. The introduction of such fields greatly enlarges the number of viable particle physics models. One of the most important properties of a model, is its symmetry: A symmetry is a structure-preserving transformation of objects such as equations, Lagrangians and action integrals. Symmetries may imply conservation laws, relate apparently distinct models, reduce the number of free parameters, protect parameters from large quantum corrections, ensure parameter relations are stable under quantum corrections and stabilize dark matter candidates against decay.
For multi-Higgs-doublet models (NHDMs), symmetry properties have been studied extensively. The simplest of such models, the two-Higgs-doublet model (2HDM) has received particular attention. The model may facilitate $\operatorname{CP}$ violation [13], necessary for generating the observed matter-antimatter asymmetry during baryogenesis [3], dark matter [14, 15, 7] and is also a necessary part of (minimal) supersymmetric theories [16, 17, 18].
Continuous symmetries connected to the identity, also known as Lie symmetries, may be studied in general by a method originally introduced by Norwegian mathematician Sophus Lie [19], and subsequently developed by his successors. The method, Lie symmetry analysis of partial differential equations (PDEs), can be utilized to find all infinitesimal (sometimes called ”local”) symmetries of a system of PDEs $\Delta$, which correspond to the Lie symmetry algebra $\mathfrak{g}$ of $\Delta$. This method is the original context were the concepts now known as Lie algebras and Lie groups emerged [19]. The Lie symmetry algebra $\mathfrak{g}$ of $\Delta$ determines to a large extent the maximal global symmetry group $\mathsf{G}$ of $\Delta$, since the Lie algebra of $\mathsf{G}$ must equal $\mathfrak{g}$. Lie symmetry analysis may be applied on almost any kind of system of PDEs, e.g. a system of Euler-Lagrange equations $E(\mathcal{L})=0$ for a theory given by a Lagrangian $\mathcal{L}$. When the dependent variables of $\mathcal{L}$ are fields, the Euler-Lagrange equations are also referred to as the field equations of the field theory given by $\mathcal{L}$. Particle physics models like multi-Higgs models are examples of field theories given by Lagrangians.
The field equations follow from the Lagrangian $\mathcal{L}$ via a variational principle. Consider the action $\mathcal{S}$,
$\mathcal{S}=\int_{\Omega}\mathcal{L}d^{4}x.$ (1.1)
Then, by varying the fields $y^{i}$ of $\mathcal{L}$ infinitesimally, with vanishing variations $\delta y^{i}$ on the boundary $\partial\Omega$ of $\Omega$, and by demanding stationary values of $\mathcal{S}$ to first order of the infinitesimal parameter, the result is the Euler-Lagrange equations. For a Lagrangian containing at most first-order derivatives (as in the 2HDM), the Euler-Lagrange equations read
$\frac{\partial\mathcal{L}}{\partial y^{i}}-d_{\mu}\frac{\partial\mathcal{L}}{\partial y^{i}_{,\mu}}=0,$ (1.2)
for each field (i.e. dependent variable) $y^{i}$ and derivative $y^{i}_{,\mu}\equiv\partial y^{i}/\partial x^{\mu}$ occurring in $\mathcal{L}$. Then there are three types of symmetries of the field equations (1.2): First, the strict variational symmetries (SVSs) which leave the action $\mathcal{S}$ strictly invariant, i.e.
$\Delta\mathcal{S}\equiv\hat{\mathcal{S}}-\mathcal{S}\equiv 0,$ (1.3)
where $\hat{\mathcal{S}}$ is the action transformed by the symmetry. Second, the divergence symmetries, where
$\Delta\mathcal{S}=\int_{\Omega}d_{\mu}\beta^{\mu}d^{4}x=\int_{\partial\Omega}\beta^{\mu}dF_{\mu}=0,$ (1.4)
is the integral of a total divergence $d_{\mu}\beta^{\mu}$, where $d_{\mu}\equiv d/dx^{\mu}$ is a total derivative, for non-vanishing, local functions $\beta^{\mu}$. Here $\Delta\mathcal{S}$ is, by the divergence theorem, converted to a boundary term, which vanishes under the usual boundary conditions on the fields, as shown in the last equality of (1.4). These first two types of symmetries are the variational symmetries, symmetries of the action $\mathcal{S}$. Variational symmetries usually translate to symmetries of the corresponding quantum theory, provided the measure of Feynman’s path integral is invariant as well as the action $\mathcal{S}$. Moreover, variational Lie symmetries lead to conserved currents by Noether’s theorem [20]. The third kind of symmetries are the non-variational symmetries that preserve the structure of the field equations, but do not leave the action $\mathcal{S}$ invariant.
When a symmetry only transforms free and dependent variables $(x,y)$, it is called a point symmetry. The derivatives of the dependent variables are also transformed under a point symmetry, but these transformations are dictated by the transformations of the variables $(x,y)$. This means a point symmetry cannot interchange a dependent variable with its derivative.
The possible, inequivalent (real) scalar point SVS groups of the 2HDM was determined in [21], using a formalism of gauge invariant, scalar bilinears. A new, discrete point transformation in the 2HDM which leads to parameter relations respected by quantum corrections, characteristic for variational symmetries, was identified in [22]. As suggested in [22], this transformation may be regarded as a complex, discrete symmetry. Such transformations may be a generic feature of quantum field theories, although they do not necessarily have to be interpreted as complex symmetries [23]. However, complex symmetries of real structures are not unknown in the context of symmetry analysis of differential equations, see e.g. [24] and [25] for examples of complex, discrete symmetries of real differential equations. An overview of the 2HDM may be found in [26]. The possible scalar, point SVSs of the 3HDM were classified in [27, 28] while the categorization of scalar symmetry groups in the 4HDM has recently been initiated in [29, 30]. In these studies, finite group theory was applied as a main tool. The possible scalar, point SVSs of the 2HDM and 3HDM, including custodial symmetries, were determined in [31, 32] and [33], respectively, by applying bilinear and tensor product methods. The scalar, Lie point SVSs of the general NHDM’s kinetic terms were determined in [34], including the corresponding symmetries in the custodial limit.
The purpose of this work is twofold: First, inspired by the aforementioned discovery in [22] of a new, complex 2HDM symmetry, we want to investigate the possibility of having Lie point divergence and non-variational symmetries in a 2HDM. Since all Higgs doublets carry the same quantum numbers (that is, the same isospin and hypercharge), unitary linear combinations of the original doublets, accompanied by appropriate redefinitions of parameters to keep the Lagrangian invariant, relate different parametrizations that describe the same physics. This phenomenon is known as basis freedom or reparametrization freedom. We will in our analysis choose specific bases were certain parameters are vanishing, to, as much as practically possible, prevent the same symmetry from occurring in multiple bases. Second, we would like to demonstrate how to classify all Lie point symmetries in models with many variables, parameters and reparametrization freedom, like the 2HDM. Both of these aims will be pursued by employing Lie symmetry analysis to the field equations of the general 2HDM.
Structure of the paper
Sections 2.1-2.3 review the relevant theory of Lie symmetry analysis of PDEs, while we present three new results in Sections 2.4 and 2.5, where the first result, Theorem 1, will be crucial for Section 4.3. We then give a simple, concrete example of how the theory of Sections 2.1-2.5 may be applied in Section 3. Section 3 is not necessary for reading Section 4, and may be skipped by readers familiar with Lie symmetry analysis of PDEs or readers who want to go straight to the main application in Section 4. In Section 4 we recall standard formalism of the 2HDM, and in Sections 4.2-4.3 we perform the Lie symmetry analysis of the 2HDM. A summary and outlook is given in Section 5, while a proof of a result of Section 2.5, Proposition 2.5, is delegated to Appendix A.
Conventions and notation
In this article we adopt the mathematicians’ convention for Lie algebras, which means that a matrix Lie group $\mathsf{G}$ with Lie algebra $\mathfrak{g}$ is generated by $\exp(\mathfrak{g})$, and not by $\exp(i\mathfrak{g})$ which would correspond to the physicists’ convention. Moreover, $d_{\mu}\equiv d/dx^{\mu}$ denotes the total derivative, while $D_{\mu}$ is reserved for the covariant derivative (2.52), unless the index is a multi-index $J$, see section 2.2, or a function $P$, implying a Fréchet derivative $D_{P}$. Repeated indices are implicitly summed over, while a check over the indices imply the summation convention is dispensed so that, for instance, $\eta^{\widecheck{\imath}}\partial_{\phi_{\widecheck{\imath}}}$ only consists of one term. Finally, we will use respectively $\Re$ and $\Im$ for real and imaginary parts of expressions, e.g. will $\lambda_{5}=\Re(\lambda_{5})+i\,\Im(\lambda_{5})$.
2 Lie symmetry analysis of PDEs
In this section we will give an overview of relevant Lie symmetry theory of PDEs. Most of the material in Sections 2.1-2.3 can be found in [35, 36, 24, 37, 38].
2.1 Point symmetries of systems of PDEs
Consider a $n$’th order system of PDEs given by
$\Delta_{i}(x,y,y^{(1)},\ldots,y^{(n)})=0,\quad\forall i\in\{1,\ldots,m\},$ (2.1)
with $d$ independent variables $x=(x^{0},\ldots,x^{d-1})$ (spacetime coordinates) and $q$ dependent variables $y=(y^{1},\ldots,y^{q})$ (the fields), where $y^{(k)}$ denotes derivatives of order $k$ of the dependent variables $y^{j}$ with respect to the independent variables $x^{\mu}$. We abbreviate (2.1) by $\Delta=0$ or just $\Delta$.
Then, a point symmetry $S$ of the system of PDEs (2.1) is a diffeomorphism [^1] Which means $S$ and $S^{-1}$ exist and have at least continuous first order derivatives, but are often taken to be smooth, i.e $C^{\infty}$. on the space of variables (i.e. it is a point transformation), which maps solutions of (2.1) to solutions. More precisely, this means
[…]
for an open set $U$, such that the transformed system of PDEs,
$\Delta_{i}(\hat{x},\hat{y},\hat{y}^{(1)},\ldots,\hat{y}^{(n)})=0,\quad\forall i\in\{1,\ldots,m\},$ (2.3)
holds whenever (2.1) holds. In (2.3) the action of $S$ is prolonged to the derivatives such that they are mapped to the corresponding derivatives in the transformed variables, that is
$S(\frac{d^{k}{y}^{i}}{d{x}^{\mu_{1}}\cdots d{x}^{\mu_{k}}})=\frac{d^{k}\hat{y}^{i}}{d\hat{x}^{\mu_{1}}\cdots d\hat{x}^{\mu_{k}}},$ (2.4)
to preserve the structure of the original system of PDEs (2.1). The condition for a diffeomorphism $S$ to be a symmetry of the system (2.1) may now be expressed in compact form as
$\Delta=0\Rightarrow\hat{\Delta}=0,$ (2.5)
[…]
2.2 Infinitesimal generators and their prolongations
An infinitesimal generator of a point transformation is a vector field
$X=\xi^{\mu}(x,y)\frac{\partial}{\partial x^{\mu}}+\eta^{i}(x,y)\frac{\partial}{\partial y^{i}},$ (2.6)
where $\mu$ is implicitly summed from 0 to d1d-1 and likewise $i$ is summed from $1$ to $q$. The infinitesimal generator of a point symmetry $S$ shows how an infinitesimal version of $S$ acts on the variables $z=(x,y)$,
$S(z)=\hat{z}=z+\epsilon X(z),$ (2.7)
for an infinitesimal $\epsilon$. The $k$-prolongation of $X$, $\operatorname{pr}^{(k)}X$, extends this action to derivatives up to $k$’th order,
$\operatorname{pr}^{(k)}X=X+\sum_{1\leq|J|\leq k}\eta_{J}^{i}\frac{\partial}{\partial y^{i}_{J}},$ (2.8)
where the multi-index $J=(j_{0},\ldots,j_{d-1})$ encodes the distinct derivatives, with norm
$|J|=j_{0}+\cdots+j_{d-1},$ (2.9)
where the derivatives of (2.8) are given by
$y^{i}_{J}\equiv\frac{\partial^{|J|}y^{i}}{(\partial x^{0})^{j_{0}}\cdots(\partial x^{d-1})^{j_{d-1}}},$ (2.10)
with coefficients
$\eta_{J}^{i}=D_{J}(Q^{i})+\xi^{\mu}\frac{\partial y^{i}_{J}}{\partial x^{\mu}}$ (2.11)
where the iterated total derivative
$D_{J}=(\frac{d}{dx^{0}})^{j_{0}}\cdots(\frac{d}{dx^{d-1}})^{j_{d-1}},$ (2.12)
and the characteristic
$Q^{i}=\eta^{i}-\xi^{\mu}\frac{\partial y^{i}}{\partial x^{\mu}}.$ (2.13)
We define the (infinite) prolongation
$\operatorname{pr}X\equiv\operatorname{pr}^{(\infty)}X$ (2.14)
by a formal, infinite series (2.8) with $k=\infty$. If $\xi^{\mu}=0$ for all $\mu$ the infinitesimal point transformation generator $X$ is said to be in evolutionary form, and for a general $X$ the infinitesimal generator
$X_{Q}=Q^{i}\partial_{y^{i}}$ (2.15)
is called its evolutionary representative. In case $Q$ includes derivatives, $X_{Q}$ will be a so-called generalized symmetry of the system of PDEs if $X$ is a symmetry thereof [35], but we will only consider cases where $X_{Q}$ is a point symmetry due to $\xi=0$ in (2.13). For a point transformation given by $X$ already in evolutionary form ($\xi=0$) its prolongation simply becomes
$\operatorname{pr}X=\operatorname{pr}X_{Q}=\sum_{i,J}(D_{J}Q^{i})\frac{\partial}{\partial y^{i}_{J}}=\sum_{i,J}(D_{J}\eta^{i})\frac{\partial}{\partial y^{i}_{J}}.$ (2.16)
In case of a very simple generator $X=y\partial_{y}$ and only one independent variable $x$ the 1-prolongation $\operatorname{pr}^{(1)}(X)=y\partial_{y}+y^{\prime}\partial_{y^{\prime}}$, and it hence just extends the infinitesimal transformation $y\to(1+\epsilon X)y=y+\epsilon y$ to the first order derivative, since $y^{\prime}\to(1+\epsilon\operatorname{pr}^{(1)}(X))y^{\prime}=y^{\prime}+\epsilon y^{\prime}$. If independent variables also transform, i.e. $\xi^{\mu}\neq 0$ for some $\mu$, the prolongations may become much more complicated, as testified by the formulas above.
Now, let $\mathsf{G}$ be a connected Lie group and let $\mathfrak{g}$ be its Lie algebra. Then $\mathsf{G}$ is a symmetry group [^2] Possibly a local group, i.e. only defined close to identity. of a fully regular333”Fully regular” here means the system $\Delta$ is locally solvable and have a non-vanishing Jacobi determinant. In practice, most systems are fully regular [36]. system of $m$ PDEs given by $\Delta=0$ if and only if
$(\operatorname{pr}X(\Delta_{i}))|_{\Delta=0}=0\quad\forall i\in\{1,\ldots,m\},$ (2.17)
[…]
The condition (2.17) is known as the linearized symmetry condition and the resulting, concrete equations are called the determining equations of $\mathfrak{g}$ (or, less precise, $\mathsf{G}$). Combining (2.8) and (2.17) shows that the set of determining equations yields an extensive, over-determined system of linear PDEs in the coefficients $\xi^{\mu}(x,y)$ and $\eta^{i}(x,y)$. This system may almost always be explicitly solved, and hence we can determine the symmetry algebra $\mathfrak{g}$, i.e. the infinitesimal symmetries. In this work we, most of the time, will apply the Mathematica package SYM [45] to calculate determining equations.
2.3 Symmetries of the action $\mathcal{S}$
Of particular interest are symmetries of the action integral
$\mathcal{S}=\int_{\Omega}\mathcal{L}({x},{y},\ldots,{y}^{(n)})\,dx^{0}\cdots dx^{d-1},$ (2.18)
since they (if Lie-type) generate Noether currents, and usually lifts to the quantized theory. As mentioned in Section 1, a transformation is called a variational symmetry of a theory, if the transformation leaves the action integral of the theory invariant,
$\mathcal{S}=\hat{\mathcal{S}}\equiv\int_{\hat{\Omega}}\mathcal{L}(\hat{x},\hat{y},\ldots,\hat{y}^{(n)})\,d\hat{x}^{0}\cdots d\hat{x}^{d-1}.$ (2.19)
A Lie-type point transformation $\exp(\epsilon X)$ for $\epsilon\in\mathbb{R}$ will be a variational symmetry if and only if
$\operatorname{pr}X(\mathcal{L})+\mathcal{L}\,{d_{\mu}\xi^{\mu}}={d_{\mu}\beta^{\mu}},$ (2.20)
where $\beta^{\mu}$ is a local [^4] The $\beta$’s will be functions on a jet space with formal coordinates $(x,y,\ldots,y^{(n)})\in\mathbb{R}^{m}$ for some $n$ and $m$, hence non-local expressions like $\int ydx$ cannot occur in the $\beta$’s [35]. function of $x$, $y$ and derivatives of $y$ up to some order, and where $d_{\mu}\equiv{d}/{dx^{\mu}}$ is a total derivative. If we can choose $\beta^{\mu}=0$ for all $\mu$, and hence
$\operatorname{pr}X(\mathcal{L})+\mathcal{L}\,{d_{\mu}\xi^{\mu}}=0,$ (2.21)
the Lie point symmetry will be a strict variational symmetry, and if we cannot choose $\beta^{\mu}=0$ for all $\mu$, it will be a divergence symmetry. For an infinitesimal Lie point divergence symmetry the action, due to the divergence theorem, transforms as
$\hat{\mathcal{S}}=\mathcal{S}+\epsilon\int_{\partial\Omega}\beta^{\mu}dF_{\mu}+\mathcal{O}(\epsilon^{2}),$ (2.22)
for a small parameter $\epsilon$, where $dF_{\mu}$ are the components of the differential outward normal vector of the boundary $\partial\Omega$ of $\Omega$. Then $\mathcal{S}$ is invariant if and only if the boundary term vanishes,
$\int_{\partial\Omega}\beta^{\mu}dF_{\mu}=0,$ (2.23)
which will be the case if the fields vanish sufficiently fast at infinity or are obeying certain periodic properties. Note that this boundary term does not depend on variations of the fields, so it is not automatically set to zero, unlike the variations at the boundary. Furthermore, it could be tempting to apply the criterion (2.20) directly to find all divergence symmetries, but it typically will lead us to very difficult, if not unsolvable, differential equations, and we are advised to proceed via the Euler-Lagrange equations, see e.g. [46].
A Lagrangian $\mathcal{L}^{\prime}=\mathcal{L}+\epsilon d_{\mu}\beta^{\mu}$ will yield the same Euler-Lagrange equations as $\mathcal{L}$, since a total divergence (also denoted a null Lagrangian) always will produce trivial Euler-Lagrange equations. In this article, the system of PDEs $\Delta=0$ will be a set of Euler-Lagrange equations of a Lagrangian $\mathcal{L}$. We define the Euler operator $E=(E_{1},\ldots,E_{q})$ with components given by formal, infinite series
$E_{i}=\sum_{|J|\geq 0}(-1)^{|J|}D_{J}\frac{\partial}{\partial y^{i}_{J}}=\frac{\partial}{\partial y^{i}}-d_{\mu}\frac{\partial}{\partial y^{i}_{,\mu}}+\ldots,\quad\;i\in\{1,\ldots,q\}$ (2.24)
with $|J|$, $y^{i}_{J}$ and $D_{J}$ given by (2.9), (2.10) and (2.12), respectively. Then, the Euler-Lagrange equations of a Lagrangian $\mathcal{L}$ may be written
$E(\mathcal{L})=0.$ (2.25)
A direct calculation show that each $E_{i}$ annihilate any total divergence $d_{\mu}\beta^{\mu}$, as implied above. In fact, it can be shown that a function $f$ is a total divergence if and only if it is annihilated by the Euler operator [35],
$E(f)=0.$ (2.26)
Therefore, if
$\operatorname{pr}X(\mathcal{L})+\mathcal{L}\,d_{\mu}\xi^{\mu}=f,$ (2.27)
for a non-vanishing $f$, the infinitesimal generator $X$ generates a divergence symmetry if and only if (2.26) holds and hence $f=d_{\mu}\beta^{\mu}$ for some local functions $\beta^{\mu}$.
The strict variational Lie point symmetries, given by the symmetry algebra $\mathfrak{g}_{\text{svar}}$, are obviously variational. The general variational Lie point symmetries, including divergence symmetries, will also generate a Lie symmetry algebra, which we denote $\mathfrak{g}_{\text{var}}$. Finally, all Lie point symmetries of the Euler-Lagrange equations of $\mathcal{L}$ will generate a Lie symmetry algebra $\mathfrak{g}_{\text{EL}}$, which will contain the two other algebras, since it can be shown that all variational symmetries are symmetries of the Euler-Lagrange equations of the theory (the converse does not hold in general). Hence, the following symmetry algebra inclusions always hold,
$\mathfrak{g}_{\text{svar}}\subseteq\mathfrak{g}_{\text{var}}\subseteq\mathfrak{g}_{\text{EL}}\,.$ (2.28)
Symmetries of $\mathfrak{g}_{\text{EL}}$ that are not included in the subalgebra $\mathfrak{g}_{\text{var}}$ are the non-variational Lie point symmetries. In section 4 we will demonstrate that for the 2HDM, the inclusions in (2.28) actually are equalities.
Due to Noether’s theorem, variational Lie point symmetries generate conserved currents. In case of a first-order Lagrangian $\mathcal{L}({x},{y},{y}^{(1)})$ a symmetry given by (2.6) induces a conservation law $d_{\mu}j^{\mu}=0$ with a conserved current
$j^{\mu}=(\eta^{i}-\xi^{\nu}y_{,\nu}^{i})\frac{\partial\mathcal{L}}{\partial y^{i}_{,\mu}}+\xi^{\mu}\mathcal{L}-\beta^{\mu}.$ (2.29)
The quantized version of the theory defined by the classical action $\mathcal{S}$, is determined by Feynman’s path integral $Z$,
$Z\;=\;\int\mathcal{D}y\;e^{\,i\mathcal{S}[y]}.$ (2.30)
Variational symmetries leaves the action invariant, and if the measure $\mathcal{D}y\equiv\prod_{i,x}dy^{i}(x)$ of $Z$ also is invariant under the transformation, the symmetry will be a symmetry of the quantized theory as well.
2.4 Theories with potentials
One of the goals of this work is to demonstrate how Lie symmetry analysis of PDEs can be applied to the Euler-Lagrange equations of particle physics models with potentials, such as multi-Higgs models. In this section we will show three results which may simplify the symmetry analysis of such models.
We start by defining the Fréchet derivative [35] of an $r$-tuple $P[y]\equiv P(x,y,\ldots,y^{(n)})$ of differential functions to be the operator $D_{P}$ which maps any $q$-tuple of differential (smooth) functions $U$ to
$D_{P}[U]\;=\;\left.\frac{d}{d\varepsilon}\,P\big[y+\varepsilon\,U\big]\right|_{\varepsilon=0}\,.$ (2.31)
Then the Fréchet derivative of a $r$-tuple of functions $P=(P_{1},\ldots,P_{r})$ is the $q\times r$ matrix with elements [35]
$(D_{P})_{ij}=\sum_{|J|\geq 0}\frac{\partial P_{i}}{\partial y_{J}^{j}}D_{J}.$ (2.32)
The adjoint $D_{P}^{\ast}$ of the Fréchet derivative then is an operator with a similar property [35]
$(D_{P})_{ij}^{\ast}=\sum_{|J|\geq 0}((-1)^{|J|}D_{J})\cdot\frac{\partial P_{j}}{\partial y_{J}^{i}}\,$, (2.33)
where e.g. $D_{x}\cdot u(x)=u_{x}+uD_{x}$ on the right side of equation (2.33). Now, let $X$ be an infinitesimal generator with coefficients
$\xi^{\mu}$ $=0,\quad\forall\mu\in\{0,\ldots,d-1\}$ (2.34)
$\eta^{i}$ $=\eta^{i}(y^{1},\ldots,y^{q}),\quad\forall i\in\{1,\ldots,q\}$ (2.35)
where the $\eta^{i}$’s are polynomials in the fields (i.e. the dependent variables). Then $X$ has characteristic
$Q^{i}=\eta^{i}(y^{1},\ldots,y^{q}),$ (2.36)
cf. (2.13), and $X$ is in evolutionary form because of (2.34). The adjoint of the Fréchet derivative of $Q$ then becomes
$(D_{Q})_{ij}^{\ast}=\frac{\partial\eta^{j}(y^{1},\ldots,y^{q})}{\partial y^{i}}=\mathcal{J}_{ji}=\mathcal{J}^{T}_{ij},$ (2.37)
where $\mathcal{J}$ is the Jacobian matrix of $\eta$, since only the $|J|=0$ term contributes in (2.33). Derivatives like $d_{\mu}$ and $\partial_{y^{i}_{,\mu}}$ do not commute. However, the following, useful commutation formula holds [35]
$E(\operatorname{pr}X_{Q}(\mathcal{L}))=\operatorname{pr}X_{Q}(E(\mathcal{L}))+D_{Q}^{\ast}E(\mathcal{L}),$ (2.38)
for a $q$-tuple of general, smooth functions $Q^{i}$, cf. (2.15). Moreover, let
$\mathcal{L}=T-V$ (2.39)
be the Lagrangian of a theory where $T$ denotes the kinetic part (the sum of the kinetic terms), while $V(\varphi_{1},\ldots,\varphi_{m})$ is a potential, that is, a real polynomial where the variables $\varphi_{j}$ form a subset of the dependent variables of the theory,
$\varphi=\{\varphi_{1},\ldots,\varphi_{m}\}\subset y=\{y^{1},\ldots,y^{q}\},$ (2.40)
where we for simplicity will (re-)order the $\varphi$’s first in $y$,
$\varphi_{i}\equiv y^{i},\quad\forall i\in\{1,2,\ldots,m\}.$ (2.41)
Note that $\varphi$ may consist of any of the fields of $y$, although the intention is to let $V(\varphi)=V(\phi)$, i.e. let $V$ be a scalar potential. Also note that the ”kinetic” part $T$ here actually may consist of any terms complementary to $V$. Moreover, we assume $E(L)=0$ does not imply any polynomial consequences of the form
$p(\varphi)=0,$ (2.42)
where $p$ is a non-trivial, multivariable polynomial in the same fields as the potential VV. [^5] In differential algebra language, $I(E(\mathcal{L}))\cap\mathbb{R}[\varphi_{1},\ldots,\varphi_{m}]=\{0\}$, where $I(E(\mathcal{L}))$ is the differential ideal generated by the Euler-Lagrange expressions $E(\mathcal{L})$ (it encodes all formal consequences of the Euler-Lagrange-equations), see e.g. [47], and $\mathbb{R}[\varphi_{1},\ldots,\varphi_{m}]$ is the ring of all polynomials in the variables $\varphi$ with real coefficients. This means we cannot derive any relation of the form (2.42), from the Euler-Lagrange equations. We will call such a theory a polynomial potential theory. Practically any multi-Higgs model will be an example of a polynomial potential theory, including the 2HDM. One reason is that each of the Euler-Lagrange equations $E_{i}(\mathcal{L})=0$ for multi-Higgs models will include distinct, non-mixed order two derivatives of the field $y^{i}$ that do not occur in any other $E_{j}(\mathcal{L})=0$ for $j\neq i$, and hence cannot be eliminated to produce some polynomial consequence (2.42). We here assume the multivariable polynomial $V$ may include any terms of any degree in the fields $\varphi$, but in some cases we will regard potentials without linear terms (but a constant in $V$ is allowed). However, terms linear in the fields will in these cases not be forbidden among the kinetic terms. We will now prove the following, useful result about evolutionary symmetries of polynomial potential theories:
Theorem 1.
Let $\mathcal{L}=T-V$ be a polynomial potential theory and let the infinitesimal generator
$X=\eta^{i}(y^{1},\ldots,y^{q})\partial_{y^{i}},$
where each $\eta^{i}$ is a polynomial, be a symmetry of $E(\mathcal{L})=0$. Moreover, assume that either $V(\varphi_{1},\ldots,\varphi_{m})$ does not contain any linear terms $\alpha_{i}\varphi_{i}$, or that $\alpha_{i}\neq 0\Rightarrow\eta^{i}(y^{1},\ldots,y^{q})$ does not contain a constant term. Then,
  1. (i)
    If $\operatorname{pr}X(T)=0$, the symmetry generated by $X$ is strictly variational.
  2. (ii)
    If $\operatorname{pr}X(T)=d_{\mu}\beta^{\mu}$ for some non-vanishing current $\beta^{\mu}$, i.e. a total divergence, the symmetry generated by $X$ is a divergence symmetry.
Proof.
The linearized symmetry condition (2.17) applied on the Euler-Lagrange equations $E(\mathcal{L})=0$ gives
$\big(\operatorname{pr}X(E_{i}(\mathcal{L}))\big)|_{E(\mathcal{L})=0}=0,\quad\forall i\in\{1,\ldots,q\},$ (2.43)
while (2.38) infer
$\operatorname{pr}X(E(\mathcal{L}))=E(\operatorname{pr}X(\mathcal{L}))-D_{Q}^{\ast}E(\mathcal{L}),$ (2.44)
since $\operatorname{pr}X=\operatorname{pr}X_{Q}$, because the former already is in evolutionary form. Assume $\operatorname{pr}X(T)=d_{\mu}\beta^{\mu}$ for a possibly vanishing current $\beta$. Substituting (2.37) in (2.44) then yields
$\operatorname{pr}X(E(\mathcal{L}))=-E(\operatorname{pr}X(V))-\mathcal{J}^{T}E(\mathcal{L}),$ (2.45)
since $E(d_{\mu}\beta^{\mu})=0$ cf. (2.26), and where $\mathcal{J}$ was the Jacobian matrix of $\eta$. Then, if $E(\mathcal{L})=0$, the first and last terms in (2.45) vanish, cf. (2.43), i.e. for any $i$
$E_{i}(\operatorname{pr}X(V))=\frac{\partial}{\partial y^{i}}\operatorname{pr}X(V)=0,$ (2.46)
since there are no derivatives in the potential. The polynomial equation (2.46) will also hold when $E(\mathcal{L})\neq 0$, since $E(\mathcal{L})=0$ does not imply any polynomial relations (i.e. consequences) between the fields of $V$, as $\mathcal{L}$ was a polynomial potential theory. But then $\operatorname{pr}X(V)=C$ for a constant $C$, and $C=0$ since all terms of
$\operatorname{pr}X(V)=X(V)=\eta^{i}\partial_{y^{i}}V$ (2.47)
are, if non-vanishing, at least linear in the fields, because if $\partial_{y^{i}}V$ includes a constant term for an $i$, then $\eta^{i}$ does not, per assumption. This means that the potential $V$ is annihilated by the prolongation of $X$,
$\operatorname{pr}X(V)=0,$ (2.48)
and hence
$\operatorname{pr}X(\mathcal{L})=\operatorname{pr}X(T)=d_{\mu}\beta^{\mu}.$ (2.49)
Thus, if $\beta=0$ then $X$ generates a strict variational symmetry, cf. (2.21) with $\xi=0$. Furthermore, if $\beta$ is non-vanishing, $X$ generates a divergence symmetry, cf. (2.27) with $\xi=0$. ∎
We will apply Theorem 1 in section 4 to show that symmetries have to be strictly variational, without having to consider all the numerous, different conditions on the parameters of the potential.
2.5 Scalar, variational symmetries in any spacetime dimension
It turns out that the purely scalar, variational Lie point symmetries of a multi-Higgs model with any number of doublets and singlets are the same for any spacetime dimension $d$. This may imply computational advantages, because if we are only interested in scalar, variational symmetries we may reduce the number of spacetime variables and gauge fields, and hence reduce the computational cost of finding the determining equations of the scalar, variational symmetries.
Let the most general NHDM+KS Lagrangian in $d$ spacetime dimensions be given by
$\mathcal{L}_{d}(x_{0},\ldots,x_{d-1})$ $=\sum_{\mu=0}^{d-1}\left(\sum_{n=1}^{N}(D_{\mu}\Phi_{n})^{\dagger}D^{\mu}\Phi_{n}+\sum_{m=1}^{K}\frac{1}{2}\partial_{\mu}s_{m}\partial^{\mu}s_{m}\right)-V(\Phi,s)+T_{\text{GB}}$ (2.50)
with $\Phi\equiv(\Phi_{1},\ldots,\Phi_{N})$ and $s\equiv(s_{1},\ldots,s_{K})$, and where the kinetic terms of the gauge bosons read
$T_{\text{GB}}=-\sum_{\mu,\nu=0}^{d-1}\left(\tfrac{1}{4}W^{a}_{\mu\nu}W^{a\mu\nu}+\tfrac{1}{4}B_{\mu\nu}B^{\mu\nu}\right).$ (2.51)
Covariant derivatives, the gauge boson field strength tensors and the Higgs doublets are given by
$D_{\mu}$ $=\partial_{\mu}+ig\frac{\sigma^{a}}{2}W^{a}_{\mu}+ig^{\prime}YB_{\mu},$ (2.52)
$W^{a}_{\mu\nu}$ $=\partial_{\mu}W^{a}_{\nu}-\partial_{\nu}W^{a}_{\mu}+g\epsilon^{abc}W^{b}_{\mu}W^{c}_{\nu},$ (2.53)
$B_{\mu\nu}$ $=\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu},$ (2.54)
$\Phi_{j}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\phi_{4(j-1)+1}+i\phi_{4(j-1)+2}\\ \phi_{4(j-1)+3}+i\phi_{4(j-1)+4}\end{pmatrix}$ (2.55)
where $\sigma^{a}$ are the Pauli matrices, $1\leq a\leq 3$. We here also include the case where $K=0$, i.e. where the theory is a pure NHDM with no gauge singlets.
Then, the variational, scalar symmetries of theories given by $\mathcal{L}_{d}$ are independent of $d$: {restatable}propositionMainProp The variational Lie point symmetries transforming only scalars are the same for all NHDM+KS Lagrangians $\mathcal{L}_{d}$, regardless of the spacetime dimension $d\in\mathbb{N}$.
Proof.
A proof is given in Appendix A. ∎
An immediate consequence of Proposition 2.5 is then
Corollary 1.
To find all scalar, variational Lie point symmetries of an NHDM+KS Lagrangian $\mathcal{L}_{4}$, it is enough to consider the simplified Lagrangian $\mathcal{L}_{1}$, since such symmetries are exactly the same for the two theories.
The simplified Lagrangian $\mathcal{L}_{1}$ will only contain 4 gauge fields and one free variable, and is hence computationally much less demanding than $\mathcal{L}_{4}$ containing 16 gauge fields and 4 free variables. The validity of Proposition 2.5 is confirmed by trying out the analysis of Section 4 on the $d=2$ variant of the 2HDM Lagrangian, $\mathcal{L}_{2}$, in addition to the $d=4$ 2HDM Lagrangian $\mathcal{L}_{4}$. [^6] The even simpler case $d=1$ generates some SYM-technical problems, since the kinetic terms of the gauge bosons vanishes in this case. However, we could have added arbitrary dummy-derivatives to circumvent this, since the kinetic gauge terms are annihilated by scalar symmetries $\operatorname{pr}X$, and hence become irrelevant for the determination of scalar symmetries. After specializing to scalar transformations, the two approaches yield exactly the same equations, and hence exactly the same symmetry algebras (all Lie symmetries of the 2HDM turn out to be strictly variational and will hence be shared by the two Lagrangians according to Proposition 2.5). However, in the $d=2$ case the number of free variables and the number of gauge fields are halved, i.e. the field equations
$E(\mathcal{L}_{2})=0$ (2.56)
form a system of 16 equations of 16 dependent and 2 free variables, while the field equations
$E(\mathcal{L}_{4})=0$ (2.57)
form a system of 24 equations of 24 dependent and 4 free variables. The computation time for SYM to calculate the determining equations was reduced to a tiny fraction in the $d=2$ case. [^7] For the 2HDM the calculation of the determining equations of $\mathcal{L}_{4}$ took nearly 14 hours on a desktop computer, while the corresponding calculation for $\mathcal{L}_{2}$ on the same computer took only 45 minutes.
In Appendix A, we argue that Proposition 2.5 and Corollary 1 cannot be extended to scalar, non-variational symmetries.
3 An illustrative example: real, scalar $\phi^{4}$ theory
The following Section may be skipped by readers familiar with Lie symmetry analysis of PDEs. The aim is here to illustrate the general theory of Sections 2.1-2.5 by a simple example, namely real, scalar $\phi^{4}$ theory.
3.1 $\phi^{4}$ theory in one dimension
Real scalar $\phi^{4}$ theory in 1d has a Lagrangian density
$\mathcal{L}=\frac{1}{2}(\phi^{\prime})^{2}-\frac{1}{2}m^{2}\phi^{2}-\frac{\lambda}{4!}\phi^{4},$ (3.1)
where $\phi=\phi(x)$ is a scalar field depending on one free variable $x$, with a non-linear field equation
$\phi^{\prime\prime}+m^{2}\phi+\frac{\lambda}{6}\phi^{3}=0.$ (3.2)
The relevant prolongation (2.8) of the infinitesimal symmetry generator $X$ of (2.6) is
$\operatorname{pr}^{(2)}X$ $=\xi(x,\phi){\partial_{x}}+\eta(x,\phi){\partial_{\phi}}+\big(\phi^{\prime}\eta_{\phi}+\eta_{x}-\phi^{\prime}\left(\phi^{\prime}\xi_{\phi}+\xi_{x}\right)\big){\partial_{\phi^{\prime}}}$
$+\Big(\phi^{\prime\prime}\eta_{\phi}+\phi^{\prime}\eta_{x\phi}+\phi^{\prime}\left(\phi^{\prime}\eta_{\phi\phi}+\eta_{x\phi}\right)+\eta_{xx}-2\phi^{\prime\prime}\left(\phi^{\prime}\xi_{\phi}+\xi_{x}\right)$
$-\phi^{\prime}\left(\phi^{\prime\prime}\xi_{\phi}+\phi^{\prime}\xi_{x\phi}+\phi^{\prime}\left(\phi^{\prime}\xi_{\phi\phi}+\xi_{x\phi}\right)+\xi_{xx}\right)\Big){\partial_{\phi^{\prime\prime}}}\,$, (3.3)
where subscripts indicate partial derivatives. We then apply (3.1) on (3.2), and thereafter use the substitution
$\phi^{\prime\prime}\to-m^{2}\phi-\frac{\lambda}{6}\phi^{3}$ (3.4)
everywhere we can, cf. the linearized symmetry condition (2.17), with the following determining equation as result,
$0$ $=\left(\tfrac{\lambda\phi^{2}}{2}+m^{2}\right)\eta-\tfrac{1}{6}\left(\lambda\phi^{3}+6m^{2}\phi\right)\left(\eta_{\phi}-3\phi^{\prime}\xi_{\phi}-2\xi_{x}\right)$
$-\phi^{\prime}\left(\phi^{\prime}\left(-\eta_{\phi\phi}+\phi^{\prime}\xi_{\phi\phi}+2\xi_{x\phi}\right)-2\eta_{x\phi}+\xi_{xx}\right)+\eta_{xx}\,.$ (3.5)
Here each coefficient of each distinct power of $\phi^{\prime}$ must vanish, for (3.1) to hold for any solution of the original field equation (3.2) (remember, $\xi$ and $\eta$ do not contain $\phi^{\prime}$), and we then get a larger system of determining equations,
${\phi^{\prime}}^{0}:\quad 0$ $=\left(\tfrac{\lambda\phi^{2}}{2}+m^{2}\right)\eta-\tfrac{1}{6}\left(\lambda\phi^{3}+6m^{2}\phi\right)\left(\eta_{\phi}-2\xi_{x}\right)+\eta_{xx}\,$,
${\phi^{\prime}}^{1}:\quad 0$ $=\tfrac{1}{2}\xi_{\phi}\left(\lambda\phi^{3}+6m^{2}\phi\right)+2\eta_{x\phi}-\xi_{xx}\,$,
${\phi^{\prime}}^{2}:\quad 0$ $=\eta_{\phi\phi}-2\xi_{x\phi}\,$,
${\phi^{\prime}}^{3}:\quad 0$ $=\xi_{\phi\phi}\,.$ (3.6)
The determining equations (3.1) is a second order system of homogeneous, linear PDEs in the coefficients of the infinitesimal generator $X$, i.e. $\xi$ and $\eta$. It can be solved by elementary methods, and the solutions will depend on whether the parameters $m^{2}$ and $\lambda$ vanishes, since vanishing parameters will yield simpler equations with more solutions.
3.1.1 Massive $\phi^{4}$ theory
Assuming
$m^{2},\lambda\neq 0$ (3.7)
the solutions of the set of four determining equations (3.1) are
$\xi(x,\phi)=c_{1},\quad\eta(x,\phi)=0,$ (3.8)
for a constant $c_{1}$, which give us the 1d Lie algebra defined by the generator
$X_{1}=\partial_{x},$ (3.9)
which generates a translational symmetry $x\to x+\epsilon$, which is abstractly just a $\mathfrak{u}(1)$ symmetry. The generator (3.9) is here a strict variational symmetry, since
$\operatorname{pr}\partial_{x}(\mathcal{L})+\mathcal{L}\,\frac{d\cdot 1}{dx}=\partial_{x}(\mathcal{L})=0,$ (3.10)
cf. (2.21), since $\mathcal{L}$ doesn’t contain $x$ explicitly.
3.1.2 Massless $\phi^{4}$ theory
If
$m=0\quad\text{and}\quad\lambda\neq 0$ (3.11)
in (3.2) and (3.1), the solutions of (3.1) are
$\xi(x,\phi)=c_{1}+c_{2}x,\quad\eta(x,\phi)=-c_{2}\phi,$ (3.12)
for arbitrary, real constants $c_{1}$ and $c_{2}$. The Lie symmetry (3.9) is hence enhanced to a 2d Lie algebra
$X_{1}$ $=\partial_{x},$
$X_{2}$ $=x\partial_{x}-\phi\partial_{\phi},$ (3.13)
where the commutator equals
$[X_{1},X_{2}]=X_{1},$ (3.14)
and we hence have the 2d, non-abelian Lie algebra $\mathfrak{a}(1)$. The generator $X_{1}$ will be strictly variational in the same manner as before. Concerning $X_{2}$,
$\operatorname{pr}^{(1)}X_{2}(\mathcal{L})+\mathcal{L}d_{x}\xi=(x\partial_{x}-\phi\partial_{\phi}-2\phi^{\prime}\partial_{\phi^{\prime}})\mathcal{L}+\mathcal{L}\,\frac{dx}{dx}=-4\mathcal{L}+\mathcal{L}=-3\mathcal{L},$ (3.15)
where $E(-3\mathcal{L})=-3E(\mathcal{L})$ is non-zero and of course proportional to the left hand side of the field equation. Hence the symmetry is non-variational, cf. (2.26), which means it is a symmetry of the field equation, but not the action. That $X_{2}$ actually is a symmetry of the field equation is confirmed by checking the linearized symmetry condition (2.17) for the field equation,
$\operatorname{pr}^{(2)}X_{2}(E(\mathcal{L}))=(x\partial_{x}-\phi\partial_{\phi}-2\phi^{\prime}\partial_{\phi^{\prime}}-3\phi^{\prime\prime}\partial_{\phi^{\prime\prime}})E(\mathcal{L})=-3E(\mathcal{L}),$ (3.16)
which vanishes when $E(\mathcal{\mathcal{L}})=0$, i.e. when the field equation (3.2) (with $m=0$) holds, and $X_{2}$ is hence a symmetry of the equation. In 4d, the corresponding symmetry will be strictly variational, cf. (3.70).
3.1.3 Free, massive scalar theory
Now, if we let
$\lambda=0\quad\text{and}\quad m\neq 0$ (3.17)
in (3.2) and (3.1), the solutions of (3.1) are
$\xi(x,\phi)$ $=d_{1}(x)+\phi d_{2}(x),$
$\eta(x,\phi)$ $=d_{3}(x)+\phi d_{4}(x)+\phi^{2}d_{2}^{\prime}(x),$ (3.18)
where
$d_{1}(x)$ $=c_{1}+\frac{c_{2}\sin(2mx)}{2m}-\frac{c_{4}\cos(2mx)}{2m^{2}}+\frac{c_{4}}{2m^{2}},$
$d_{2}(x)$ $=c_{5}\cos(mx)+c_{6}\sin(mx),$
$d_{3}(x)$ $=c_{7}\cos(mx)+c_{8}\sin(mx),$
$d_{4}(x)$ $=-c_{2}\sin^{2}(mx)+c_{3}+\frac{c_{4}\sin(2mx)}{2m}.$ (3.19)
Then the infinitesimal generator (2.6) is given by
$X=\sum_{i=1}^{8}c_{i}X_{i},$ (3.20)
where the generators $X_{i}$ may be taken as
$X_{1}$ $=(1/m)\partial_{x},$
$X_{2}$ $={(1/m)\sin(mx)\cos(mx){\partial_{x}}}-\sin^{2}(mx)\phi{\partial_{\phi}},$
$X_{3}$ $=\phi{\partial_{\phi}},$
$X_{4}$ $=(1/m)\sin^{2}(mx){\partial_{x}}+\sin(mx)\cos(mx)\phi{\partial_{\phi}},$
$X_{5}$ $=(1/m)\cos(mx)\phi{\partial_{x}}-\sin(mx)\phi^{2}{\partial_{\phi}},$
$X_{6}$ $=(1/m)\sin(mx)\phi{\partial_{x}}+\cos(mx)\phi^{2}{\partial_{\phi}},$
$X_{7}$ $=\cos(mx){\partial_{\phi}},$
$X_{8}$ $=\sin(mx)\partial_{\phi},$ (3.21)
where the factors $(1/m)$ ensure we will get no occurences of the mass parameter $m$ in the commutator table. Here $X_{1}$ corresponds to translations and $X_{3}$ scaling symmetries. Some occurrences of the parameter $m^{-1}$ have here been absorbed in the constants $c_{j}$, which should be kept in mind if we want to explore the symmetries corresponding to the case $m=0$. The commutator table of (3.1.3) may then be calculated by applying SYM [45]
$\small{\begin{array}[]{|c|c|c|c|c|c|c|c|c|}\hline\cr\text{}&X_{1}&X_{2}&X_{3}&X_{4}&X_{5}&X_{6}&X_{7}&X_{8}\\ \hline\cr X_{1}&0&X_{1}-2X_{4}&0&2X_{2}+X_{3}&-X_{6}&X_{5}&-X_{8}&X_{7}\\ \hline\cr X_{2}&2X_{4}-X_{1}&0&0&X_{4}&-X_{5}&0&0&X_{8}\\ \hline\cr X_{3}&0&0&0&0&X_{5}&X_{6}&-X_{7}&-X_{8}\\ \hline\cr X_{4}&-2X_{2}-X_{3}&-X_{4}&0&0&-X_{6}&0&-X_{8}&0\\ \hline\cr X_{5}&X_{6}&X_{5}&-X_{5}&X_{6}&0&0&X_{4}-X_{1}&X_{3}-X_{2}\\ \hline\cr X_{6}&-X_{5}&0&-X_{6}&0&0&0&-X_{2}-2X_{3}&-X_{4}\\ \hline\cr X_{7}&X_{8}&0&X_{7}&X_{8}&X_{1}-X_{4}&X_{2}+2X_{3}&0&0\\ \hline\cr X_{8}&-X_{7}&-X_{8}&X_{8}&0&X_{2}-X_{3}&X_{4}&0&0\\ \hline\cr\end{array}}$ (3.31)
We note that $\{X_{1},X_{7},X_{8}\}$, $\{X_{1},X_{5},X_{6}\}$ and $\{X_{1},X_{3},X_{7},X_{8}\}$ are subalgebras. The subalgebra $\{X_{3},X_{6}\}=\mathfrak{a}(1)$ is non-compact, so then the ambient algebra is non-compact as well. We will show that the Lie algebra generated by (3.1.3) is $\mathfrak{sl}(3)$ in Section (3.1.4). The only scalar symmetry $X_{3}$ has a prolongation which acts as
$\operatorname{pr}^{(1)}X_{3}(\mathcal{L})=(\phi\partial_{\phi}+\phi^{\prime}\partial_{\phi^{\prime}})(\mathcal{L})=2\mathcal{L},$ (3.32)
and is hence not a variational symmetry, since the Euler operator of the result of (3.32) evidently is not zero: $E(2\mathcal{L})=2E(\mathcal{L})$ is a multiple of the Euler-Lagrange-expression, which is not identically zero. Of course, the expression vanishes when assuming the Euler-Lagrange equation hold, but we don’t assume this on the level of variational (also called ”off-shell”) symmetries.
3.1.4 Free, massless scalar theory
In case
$m=\lambda=0$ (3.33)
the solutions of the determining equations (3.1) still yield a Lie algebra that is 8d, with generators
$X_{1}$ $=\partial_{x},$
$X_{2}$ $=x\partial_{x},$
$X_{3}$ $=x^{2}\partial_{x}+x\phi\partial_{\phi},$
$X_{4}$ $=\phi\partial_{x},$
$X_{5}$ $=x\phi\partial_{x}+\phi^{2}\partial_{\phi},$
$X_{6}$ $=\partial_{\phi},$
$X_{7}$ $=\phi\partial_{\phi},$
$X_{8}$ $=x\partial_{\phi},$ (3.34)
and commutator table
$\small{\begin{array}[]{|c|c|c|c|c|c|c|c|c|}\hline\cr\text{}&X_{1}&X_{2}&X_{3}&X_{4}&X_{5}&X_{6}&X_{7}&X_{8}\\ \hline\cr X_{1}&0&X_{1}&2X_{2}+X_{7}&0&X_{4}&0&0&X_{6}\\ \hline\cr X_{2}&-X_{1}&0&X_{3}&-X_{4}&0&0&0&X_{8}\\ \hline\cr X_{3}&-2X_{2}-X_{7}&-X_{3}&0&-X_{5}&0&-X_{8}&0&0\\ \hline\cr X_{4}&0&X_{4}&X_{5}&0&0&-X_{1}&-X_{4}&X_{7}-X_{2}\\ \hline\cr X_{5}&-X_{4}&0&0&0&0&-X_{2}-2X_{7}&-X_{5}&-X_{3}\\ \hline\cr X_{6}&0&0&X_{8}&X_{1}&X_{2}+2X_{7}&0&X_{6}&0\\ \hline\cr X_{7}&0&0&0&X_{4}&X_{5}&-X_{6}&0&-X_{8}\\ \hline\cr X_{8}&-X_{6}&-X_{8}&0&X_{2}-X_{7}&X_{3}&0&X_{8}&0\\ \hline\cr\end{array}}$ (3.44)
We again note that the algebra has an $\mathfrak{a}(1)=\{X_{1},X_{2}\}$ subalgebra, so the 8d algebra cannot be a compact algebra, since the subalgebra $\mathfrak{a}(1)$ is not compact. In fact, this algebra is $\mathfrak{sl}(3)$ (cf. Olver’s no 6.8 basis e.g. in [43]).
Lie himself showed that a second order equation with an 8d algebra, is always equivalent to the massless equation $\phi^{\prime\prime}=0$. The equation (3.2) with $\lambda=0,m^{2}\neq 0$ is equivalent to the massless equation $\hat{\phi}^{\prime\prime}=0$ through the point transformation (see e.g. [48])
$\hat{x}=\tan(mx),\quad\hat{\phi}=\frac{\phi}{\cos(mx)},$ (3.45)
which makes the massless equation hold if and only if the massive equation holds, as long as $\cos(mx)\neq 0$. Note that the transformation (3.45) is not a symmetry, since the structure of the equation is not conserved.
Moreover, the algebra corresponding to (3.1.3) and (3.31) is $\mathfrak{sl}(3)$, just as for the case $m=0$. First, this is the only possible 8d algebra for any second order ODE. Second, the generators (3.1.3) may be transformed to generators yielding the exactly same commutator table as for $m=0$ (3.44), by a basis shift:
By applying the transformations (3.45) and their inverses
$x=\frac{\arctan(\hat{x})}{m},\quad\phi=\frac{\hat{\phi}}{\sqrt{1+\hat{x}^{2}}},$ (3.46)
in combination with chain rules
$\partial_{x}$ $=(\partial_{x}\hat{x})\partial_{\hat{x}}+(\partial_{x}\hat{\phi})\partial_{\hat{\phi}}$,
$\partial_{\phi}$ $=(\partial_{\phi}\hat{x})\partial_{\hat{x}}+(\partial_{\phi}\hat{\phi})\partial_{\hat{\phi}}$, (3.47)
we get a correspondence between the generators (3.1.3) related to $m\neq 0$, and the generators (3.1.4) corresponding to $m=0$. We will here denote the latter as hatted quantities, consistent with the transformation (3.45). For instance,
[…]
$=\partial_{\hat{x}}+\hat{x}^{2}\partial_{\hat{x}}+\hat{x}\hat{\phi}\partial_{\hat{\phi}}$
$=\hat{X}_{1}+\hat{X}_{3},$ (3.48)
where the last line refers to the generators of (3.1.4). Then the generators of (3.1.3) and (3.1.4) are connected through
$X=T\hat{X}$, (3.49)
where the matrix $T$ equals
$T=\left(\begin{array}[]{cccccccc}1&0&1&0&0&0&0&0\\ 0&1&0&0&0&0&0&0\\ 0&0&0&0&0&0&1&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1\\ \end{array}\right),$ (3.58)
and we can express the generators (3.1.3) in a new basis given by
${X}^{\prime}=T^{-1}X.$ (3.59)
Then a calculation of the commutation table in the basis ${X}^{\prime}$, shows that the table now is identical to the one in (3.44), the $m=0$ case. Hence we have shown explicitly how the Lie symmetry algebra of the free, massive equation is the same as for the corresponding massless equation, namely $\mathfrak{sl}(3,\mathbb{R})$.
As in the massive case, the generator $X_{7}=\phi\partial_{\phi}$ will correspond to a non-variational symmetry, since its prolongation again will map $\mathcal{L}$ to a multiple of itself. The new scalar symmetry $X_{6}=\partial_{\phi}$ equals its own prolongation and will annihilate $\mathcal{L}=(1/2)(\phi^{\prime})^{2}$, and is thus a strict variational symmetry. The accordance with Theorem 1 is here trivial, since $V=0$.
3.2 $\phi^{4}$ theory in (3+1)d spacetime
Real, scalar $\phi^{4}$ theory in 4d has a Lagrangian density
$\mathcal{L}=\frac{1}{2}\partial_{\mu}\phi\,\partial^{\mu}\phi-\frac{1}{2}m^{2}\phi^{2}-\frac{\lambda}{4!}\phi^{4},$ (3.60)
with a non-linear field equation
$\partial_{\mu}\partial^{\mu}\phi+m^{2}\phi+\frac{\lambda}{6}\phi^{3}=0,$ (3.61)
We may now analyze the symmetries of (3.61), e.g. by applying MathLie [49] or SYM [45]. The single equation (3.61) here yields 19 determining equations.
3.2.1 The case $m,\lambda\neq 0$
In case $m,\lambda\neq 0$, the Lie symmetry algebra consists of 10 generators, the 6 first corresponding to the Lorentz algebra $\mathfrak{so}(1,3)$,
$X_{1}$ $=x_{0}\partial_{x_{1}}+x_{1}\partial_{x_{0}},$
$X_{2}$ $=x_{0}\partial_{x_{2}}+x_{2}\partial_{x_{0}},$
$X_{3}$ $=x_{0}\partial_{x_{3}}+x_{3}\partial_{x_{0}},$
$X_{4}$ $=x_{2}\partial_{x_{3}}-x_{3}\partial_{x_{2}},$
$X_{5}$ $=x_{3}\partial_{x_{1}}-x_{1}\partial_{x_{3}},$
$X_{6}$ $=x_{1}\partial_{x_{2}}-x_{2}\partial_{x_{1}},$ (3.62)
where the three first generators are the boosts in $x_{1},x_{2}$ and $x_{3}$ directions, and the three last are the rotations about the $x_{1},x_{2}$ and $x_{3}$ axis, respectively. Together with the 4 translations
$X_{7}$ $=\partial_{x_{0}},$
$X_{8}$ $=\partial_{x_{1}},$
$X_{9}$ $=\partial_{x_{2}},$
$X_{10}$ $=\partial_{x_{3}},$ (3.63)
(3.2.1) yields the Poincaré algebra $\mathfrak{iso}(1,3)$, which is the symmetry algebra for the case $m,\lambda\neq 0$.
3.2.2 The case $m=0$ and $\lambda\neq 0$
In the case of $\phi^{4}$ theory in (3+1)d spacetime with $m=0$ and $\lambda\neq 0$ SYM [45] reveals the symmetry algebra $\mathfrak{iso}(1,3)$ is enhanced with the five generators [^8] Thanks to Stylianos Dimas for providing a tutorial Mathematica notebook demonstrating eqs. (3.2.1)–(3.2.2) and (3.2.3) using SYM.
$X_{11}$ $=\Sigma_{i=1}^{4}x_{i}\partial_{x_{i}}-\phi\partial_{\phi}$
$X_{12}$ $=-2\phi x_{0}\,\partial_{\phi}+2x_{0}x_{1}\,\partial_{x_{1}}+2x_{0}x_{2}\,\partial_{x_{2}}+2x_{0}x_{3}\,\partial_{x_{3}}+(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})\,\partial_{x_{0}},$
$X_{13}$ $=-2\phi x_{1}\,\partial_{\phi}+2x_{0}x_{1}\,\partial_{x_{0}}+2x_{1}x_{2}\,\partial_{x_{2}}+2x_{1}x_{3}\,\partial_{x_{3}}+(x_{0}^{2}+x_{1}^{2}-x_{2}^{2}-x_{3}^{2})\,\partial_{x_{1}},$
$X_{14}$ $=-2\phi x_{2}\,\partial_{\phi}+2x_{0}x_{2}\,\partial_{x_{0}}+2x_{1}x_{2}\,\partial_{x_{1}}+2x_{2}x_{3}\,\partial_{x_{3}}+(x_{0}^{2}-x_{1}^{2}+x_{2}^{2}-x_{3}^{2})\,\partial_{x_{2}},$
$X_{15}$ $=-2\phi x_{3}\,\partial_{\phi}+2x_{0}x_{3}\,\partial_{x_{0}}+2x_{1}x_{3}\,\partial_{x_{1}}+2x_{2}x_{3}\,\partial_{x_{2}}+(x_{0}^{2}-x_{1}^{2}-x_{2}^{2}+x_{3}^{2})\,\partial_{x_{3}}.$ (3.64)
Scaling symmetries like $X_{11}$ are known for not being carried over to the quantum theory, due to the regulator introducing a scale. We will now demonstrate that the scaling symmetry generated by $X_{11}$ is, nevertheless, strictly variational, in contrast to the situation in the 1d case. The generator $X_{11}$ corresponds to infinitesimal symmetry transformations
$x^{\mu}\to(1+\epsilon)x^{\mu},\quad\phi\to(1-\epsilon)\phi.$ (3.65)
Then the derivative transforms as,
$\frac{\partial\phi}{\partial x^{\mu}}\to\frac{(1-\epsilon)\partial\phi}{(1+\epsilon)\partial x^{\mu}}=(1-2\epsilon)\frac{\partial\phi}{\partial x^{\mu}},$ (3.66)
while the measure transforms as
$d^{4}x\to(1+4\epsilon)d^{4}x$ (3.67)
to first order in the infinitesimal $\epsilon$, cf. (3.65). Then, suppressing $\epsilon$, $\delta\phi=-\phi$, $\delta x=x$, $\delta({\partial_{\mu}\phi})=-2{\partial_{\mu}\phi}$ and $\delta(d^{4}x)=4d^{4}x$, which gives a first order variation
$\frac{\delta(\mathcal{L}\,d^{4}x)}{d^{4}x}=\frac{\partial\mathcal{L}}{\partial\phi}\delta\phi+\frac{\partial\mathcal{L}}{\partial(\partial_{\mu}\phi)}\delta({\partial_{\mu}\phi})+\frac{\mathcal{L}\delta(d^{4}x)}{d^{4}x}=\frac{\lambda\phi^{4}}{6}-2\partial_{\mu}\phi\partial^{\mu}\phi+4\mathcal{L}=0,$ (3.68)
of the Lagrangian of $\phi^{4}$ theory. We get the same result by applying the 1-prolongation of $X_{11}$,
$\operatorname{pr}^{(1)}X_{11}=x^{\mu}\partial_{\mu}-\phi\partial_{\phi}-2(\partial_{\mu}\phi)\frac{\partial}{\partial(\partial_{\mu}\phi)},$ (3.69)
to $\mathcal{L}$, since the definition (2.21) of a strict variational symmetry then holds,
$\operatorname{pr}^{(1)}X_{11}(\mathcal{L})+\mathcal{L}\,d_{\mu}\xi^{\mu}=-4\mathcal{L}+4\mathcal{L}=0,$ (3.70)
where $\xi^{\mu}=x^{\mu}$, in contrast to the 1d case (3.15). There are no scalar symmetries, just as the case was for the corresponding 1d theory of Section 3.1.2. This is in accordance with Corollary 1, which says the scalar, variational symmetries will be the same for the 1d and the 4d theories.
3.2.3 The case $m\neq 0$ and $\lambda=0$
In the case of $\phi^{4}$ theory in (3+1)d spacetime with $m\neq 0$ and $\lambda=0$ SYM [45] shows the symmetry algebra $\mathfrak{iso}(1,3)$ is enhanced with the generators
$X_{11}$ $=\phi\partial_{\phi},$
$X_{12}$ $=\mathcal{F}_{1}(x)\,\partial_{\phi},$ (3.71)
where $\mathcal{F}_{1}(x)$ is an arbitrary function of the spacetime variables, satisfying
$\partial_{\mu}\partial^{\mu}\mathcal{F}_{1}+m^{2}\mathcal{F}_{1}=0.$ (3.72)
Hence the symmetry algebra in this case becomes infinite dimensional, with $\mathfrak{iso}(1,3)$ as a finite subalgebra. The only scalar symmetry $X_{11}$ here is the same as for the 1d case in Section 3.1.3. By Corollary 1, the scalar, variational symmetries of the 4d theory must be the same as those of the 1d theory, which can easily be confirmed: As for the 1d case the prolongation of $X_{11}$ just maps $\mathcal{L}$ to a multiple of itself, and is hence again non-variational.
3.2.4 The case $m=\lambda=0$
In the case $m=\lambda=0$ the Poincaré algebra $\mathfrak{iso}(1,3)$ is enhanced with both sets of generators given by (3.2.2) and (3.2.3), and hence the algebra in this case again becomes infinite dimensional. The symmetry $X_{11}=\phi\partial_{\phi}$ will still map $\mathcal{L}$ to a multiple of itself, as in Section 3.1.4, and is hence non-variational as before. The 4d, scalar symmetry is hence of the same nature as in the 1d case, in harmony with Corollary 1.
No divergence symmetries were found in the analysis in Section 3, but they occur e.g. in scalar gauge singlet extensions of the SM [50]. We finally also note all three cases $m=0$, $\lambda=0$ and $\lambda=m=0$ may be imposed by symmetries, since all cases are corresponding to distinct symmetries.
4 2HDM
The Lagrangian of a 2HDM is defined by (2.50) with $N=2$ and $K=0$,
$\mathcal{L}=-\tfrac{1}{4}W^{a}_{\mu\nu}W^{a\mu\nu}-\tfrac{1}{4}B_{\mu\nu}B^{\mu\nu}+\sum_{i=1}^{2}(D_{\mu}\Phi_{i})^{\dagger}(D^{\mu}\Phi_{i})-V(\Phi_{1},\Phi_{2}),$ (4.1)
with Higgs doublets
$\Phi_{1}=\frac{1}{\sqrt{2}}\begin{pmatrix}\phi_{1}+i\phi_{2}\\ \phi_{3}+i\phi_{4}\end{pmatrix},\quad\Phi_{2}=\frac{1}{\sqrt{2}}\begin{pmatrix}\phi_{5}+i\phi_{6}\\ \phi_{7}+i\phi_{8}\end{pmatrix}$, (4.2)
where covariant derivatives and the gauge boson field strength tensors are given by (2.52)-(2.54). The most general, renormalizable 2HDM potential may be written as
$V(\Phi_{1},\Phi_{2})$ $=m_{11}^{2}\,\Phi_{1}^{\dagger}\Phi_{1}+m_{22}^{2}\,\Phi_{2}^{\dagger}\Phi_{2}-\big[m_{12}^{2}\,\Phi_{1}^{\dagger}\Phi_{2}+\text{h.c.}\big]$
$\quad+\tfrac{\lambda_{1}}{2}(\Phi_{1}^{\dagger}\Phi_{1})^{2}+\tfrac{\lambda_{2}}{2}(\Phi_{2}^{\dagger}\Phi_{2})^{2}+\lambda_{3}(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{2})+\lambda_{4}(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{1})$
$\quad+\Big[\tfrac{\lambda_{5}}{2}(\Phi_{1}^{\dagger}\Phi_{2})^{2}+\lambda_{6}(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{1}^{\dagger}\Phi_{2})+\lambda_{7}(\Phi_{2}^{\dagger}\Phi_{2})(\Phi_{1}^{\dagger}\Phi_{2})+\text{h.c.}\Big],$ (4.3)
where ”h.c.” means Hermitian conjugate, i.e. conjugate transpose, with complex parameters $m_{12}^{2}$, $\lambda_{5}$, $\lambda_{6}$ and $\lambda_{7}$, other parameters real. By a basis shift (i.e. a reparametrization) of the doublets,
$(\Phi_{1},\Phi_{2})^{T}\to(\hat{\Phi}_{1},\hat{\Phi}_{2})^{T}=U(\Phi_{1},\Phi_{2})^{T}$ (4.4)
with [^9] Or with $U\in\mathsf{SU}(2)$, if we absorb the complex phase of $\mathsf{U}(2)$ into $\mathsf{U}(1)_{Y}$ hypercharge symmetry. $U\in\mathsf{U}(2)$, we may diagonalize the mass-squared matrix and remove the complex phase of e.g. $\lambda_{5}$, which means
$m_{12}^{2}=0,\qquad\Im(\lambda_{5})=0,$ (4.5)
in the potential $\hat{V}(\hat{\Phi}_{1},\hat{\Phi}_{2})=V({\Phi}_{1},{\Phi}_{2})$ written in the new, hatted basis.
The 24 Euler-Lagrange equations $E(\mathcal{L})=0$ of (4.1) are of the form
$\frac{\partial\mathcal{L}}{\partial y^{i}}-d_{\mu}\frac{\partial\mathcal{L}}{\partial y^{i}_{,\mu}}=0,\quad 1\leq i\leq 24,$ (4.6)
with
$\{y^{i}\}_{i=1}^{24}=\{\phi_{1},\ldots,\phi_{8},B_{0},\ldots,B_{3},W_{0}^{1},W_{0}^{2},\ldots,W_{3}^{3}\}.$ (4.7)
These equations will be variants of interacting (i.e. nonlinear) Klein-Gordon and Proca equations.
4.1 A bilinear formalism
The 2HDM potential may also be written in a particularly useful manner by gauge invariant bilinears [51, 52].
$V=M_{0}r_{0}+M_{a}r_{a}+\Lambda_{0}r_{0}^{2}+L_{a}r_{0}r_{a}+r_{a}\Lambda_{ab}r_{b}$ (4.8)
by employing bilinears in the fields,
$r_{0}=\Phi_{i}^{\dagger}\Phi_{i}\;,\quad r_{a}=\Phi_{i}^{\dagger}(\sigma_{a})_{ij}\Phi_{j},$ (4.9)
where $a\in\{1,2,3\}$ and bilinears $r_{a}$ are given in terms of the Pauli matrices $\sigma_{a}$, which infer
$r_{1}$ $=\frac{1}{2}(\Phi_{1}^{\dagger}\Phi_{2}+\Phi_{2}^{\dagger}\Phi_{1}),\quad r_{2}=-\frac{i}{2}(\Phi_{1}^{\dagger}\Phi_{2}-\Phi_{2}^{\dagger}\Phi_{1}),$
$r_{3}$ $=\frac{1}{2}(\Phi_{1}^{\dagger}\Phi_{1}-\Phi_{2}^{\dagger}\Phi_{2}).$ (4.10)
The parameters of (4.8) are given by
$M_{0}$ $=m_{11}^{2}+m_{22}^{2},\quad\Lambda_{0}=\tfrac{1}{2}(\lambda_{1}+\lambda_{2})+\lambda_{3},$ (4.11)
$L$ $=\left(-\Re(\lambda_{6}+\lambda_{7}),\Im(\lambda_{6}+\lambda_{7}),\frac{1}{2}(\lambda_{2}-\lambda_{1})\right)^{T}$ (4.12)
$M$ $=\left(2\Re(m_{12}^{2}),-2\Im(m_{12}^{2}),m_{22}^{2}-m_{11}^{2}\right)^{T},$ (4.13)
[…]
Under a change of basis
$\Phi_{i}\rightarrow U_{ij}\Phi_{j}\,,\quad U\in\mathsf{SU}(2),$ (4.18)
$r_{0}$ is a singlet while $r_{a}$ transforms under the adjoint representation of $\mathsf{SU}(2)$,
$r_{0}\rightarrow r_{0}\,,\quad r_{a}\rightarrow R_{ab}(U)r_{b},$ (4.19)
with
$R_{ab}(U)=\frac{1}{2}\text{Tr}(U^{\dagger}\lambda_{a}U\lambda_{b}).$ (4.20)
To keep the potential invariant, the coupling constants then transforms as follows under this change of basis,
$\Lambda$ $\to R(U)\Lambda R^{T}(U),$ (4.21)
$L$ $\to R(U)L,$ (4.22)
$M$ $\to R(U)M.$ (4.23)
Now, since $\text{Ad}_{\mathsf{SU}(2)}=\mathsf{SO}(3)$ all matrices of $\mathsf{SO}(3)$ correspond to a $\mathsf{SU}(2)$ Higgs basis transformation, and we may always diagonalize the matrix $\Lambda$ of (4.17) by a basis transformation. Given a arbitrary potential, we may change basis in such a way that $\Lambda$ is diagonalized. This means that
$\Im(\lambda_{5})=0,\quad\lambda_{7}=\lambda_{6},$ (4.24)
in the new basis. The basis shift will of course depend on the values of the parameters of $\Lambda$ in the original basis, and we have hence reduced the number of parameters by three. A slightly different bilinear formalism was applied in the beautiful derivation of the strict variational symmetries of the 2HDM in [21]. A crucial part of the proof is the aforementioned diagonalizability of $\Lambda$, which does not carry over to the analogous matrix $\Lambda$ for higher $N$, since $\text{Ad}_{\mathsf{SU}(N)}\varsubsetneq\mathsf{SO}(N^{2}-1)$ is dramatically smaller when $N>2$. This again makes classification of symmetries by bilinear formalisms much harder for $N>2$.
4.2 Finding and solving determining equations
We now find the determining equations (2.17) of the system of PDEs given by (4.6), by applying SYM [45]. The system (4.6) yields a set of 1733 determining equations, all linear PDEs in the coefficients $\xi^{\mu}$ and $\eta^{i}$ of (2.6). Since we are only interested in scalar symmetries, we set
$\xi^{\mu}$ $=0,\quad\text{for all}\quad 0\leq\mu\leq 3,$
$\eta^{i}$ $=0,\quad\text{for all}\quad 9\leq i\leq 24,$ (4.25)
which means spacetime variables and gauge bosons are kept constant under the considered transformations. Then the simplest (shortest) of the remaining determining equations read
$\partial_{\phi_{j}}\partial_{\phi_{k}}\eta^{i}=0,\quad\text{for all}\quad 1\leq i,j,k\leq 8,$ (4.26)
and
$\partial_{y^{j}}\eta^{i}=0\quad\text{for all}\quad j>8\wedge 1\leq i\leq 8,$ (4.27)
which means the $\eta^{i}$’s are affine (or sometimes referred to as linear) in the scalar fields, and do not depend on the gauge fields. Hence, we may proceed solving the determining equations by substituting
$\eta^{i}=a_{i}+b_{ij}\phi_{j},$ (4.28)
with an implicit sum over $j$ from 1 to 8, into the determining equations. We don’t include explicit spacetime variables in (4.28), since we are only considering pure scalar symmetries, although it would have been a principal possibility for non-scalar symmetries. After substituting (4.28), we are left with a system of 353 polynomial equations
$P_{k}(y^{1},\ldots,y^{q})=0$ (4.29)
in the $q=24$ fields, with $1\leq k\leq 353$. The polynomials $P_{k}$ will contain parameters from the Lagrangian, in addition to the $a_{i}$’s and $b_{ij}$’s. All the equations (4.29) have to hold for all values of the fields, to have a symmetry. This means that the coefficients of all distinct monomials in the fields, in all equations (4.29), have to be zero. More precisely, let a general monomial $\bar{m}$ in the $q=24$ fields be written
$\bar{m}(n_{1},\ldots,n_{q})=(y^{1})^{n_{1}}\cdots(y^{q})^{n_{q}},\quad n_{j}\in\mathbb{N}_{0},$ (4.30)
where $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. Then
$C\bar{m}(n_{1},\ldots,n_{q})\in P_{k}\Rightarrow C=0$ (4.31)
when $C\equiv C_{n_{1},\ldots,n_{q}}(m^{2},\lambda,g,g^{\prime}Y)$ is the maximal coefficient of $\bar{m}(n_{1},\ldots,n_{q})$, that is
$D\bar{m}(n_{1},\ldots,n_{q})\notin(P_{k}-C\bar{m}(n_{1},\ldots,n_{q})),$ (4.32)
for any, non-zero coefficient $D\equiv D_{n_{1},\ldots,n_{q}}(m^{2},\lambda,g,g^{\prime}Y)$. This splitting in monomials yields 1412 equations of the form $C=0$, some of them very simple. The 80 simplest equations are of the type
$h\,a_{i}=0,\quad\text{and}\quad h\,b_{ij}=0,\quad\text{with}\quad h\in\{g,g^{\prime}Y,gg^{\prime}Y\},$ (4.33)
and may be solved with solutions
$a_{i}=0\quad\forall i\in\{1,\ldots,8\},$ (4.34)
and
$b_{ij}=0$ (4.35)
for all $i,j$ except for indices corresponding to any of the following 24 parameters
$B_{1}=\{$ $b_{12},\,b_{15},\,b_{16},\,b_{21},\,b_{25},\,b_{26},\,b_{34},\,b_{37},\,b_{38},\,b_{43},\,b_{47},\,b_{48},$
$b_{51},\,b_{52},\,b_{56},\,b_{61},\,b_{62},\,b_{65},\,b_{73},\,b_{74},\,b_{78},\,b_{83},\,b_{84},\,b_{87}\},$ (4.36)
parameters which do not vanish at this stage. We also could have considered the case $g^{\prime}=0$ here, and hence investigated custodial symmetries [53, 54, 34, 32, 55, 56, 57, 58], but we will refrain from it, as these of course are not exact symmetries of the considered Lagrangian (4.1), and as this could have doubled up the analysis. Eq. (4.34) means there cannot be any pure scalar shift symmetries of the field equations in a 2HDM. We continue our process of solving the determining equations by substituting the solutions (4.34) and (4.35) into the 1412 determining equations, where only the parameters (4.2) are kept non-zero among the $b_{ij}$’s.
At this stage we also implement (4.5), i.e. set $m_{12}^{2}=0$ and $\Im(\lambda_{5})=0$, although we just as well could have done it from the start, before calculating the Euler-Lagrange equations. We then end up with a system
$\mathcal{D}_{i}=0$ (4.37)
of 548 determining equations.
4.3 Parameter cases and reductions of the potential
Let a reduced potential be a potential where one or more parameters have been eliminated, by transforming the potential to a specific scalar basis, through an $\mathsf{U}(2)$ Higgs basis transformation (4.4). Reduced potentials have been discussed and employed in e.g. [13, 59, 21, 32]. We proceed by considering different parameter cases, and will try to keep the potential optimally reduced in each case. This means that we eliminate as many parameters as possible through an appropriate basis transformation, and hence, more or less, lock the basis and thereby discard reparametrization freedom. We will do this to avoid that the same symmetry may manifest itself in different bases of the doublets excessively, typically by Lie algebras with generators involving potential parameters. Although all these manifestations of the same symmetry will be detected by the Lie symmetry analysis (for a clarification of this, see Appendix B), we want to avoid this as much as possible, because we otherwise would have to show that two different manifestations of the same symmetry in fact are equivalent, which may imply extra work.
4.3.1 The case $m_{11}^{2}\neq m_{22}^{2}$
In addition to the diagonalization of the mass-squared matrix, and hence the reduction (4.5) of the potential, we now assume
$m_{11}^{2}\neq m_{22}^{2}$ (4.38)
Then the 164 simplest determining equations are of the forms
$(m_{11}^{2}-m_{22}^{2})b_{ij}$ $=0\Rightarrow b_{ij}=0$ (4.39)
$h(b_{ij}\pm b_{kl})$ $=0\Rightarrow b_{ij}=\mp b_{kl}$ (4.40)
for $h>0$ cf. (4.33) and certain indices $i,j,k$ and $l$. Solving these equations gives
$-b_{12}$ $=-b_{34}=b_{43}=b_{21},$
$b_{56}$ $=-b_{65}=-b_{87}=b_{78},$
$b_{ij}$ $=0,\quad\text{for other}\;\>i,j,$ (4.41)
which means we have only two free parameters $b_{21}$ and $b_{78}$. Inserting (4.3.1) into the full set of determining equations, and then applying Mathematica’s built-in Reduce function, gives us that either
$b_{21}+b_{78}=0$ (4.42)
or
$\Re(\lambda_{5})=\Re(\lambda_{6})=\Re(\lambda_{7})=\Im(\lambda_{6})=\Im(\lambda_{7})=0$ (4.43)
Here, the case (4.42) substituted into (4.28) and then into the infinitesimal generator (2.6) results in a 1-dimensional algebra that is present for all [^10] This corresponds to the fact there are no restricitions on the parameters of the potential in (4.42). potentials, given by the generator
$X_{Y}=-\phi_{2}\partial_{\phi_{1}}+\phi_{1}\partial_{\phi_{2}}-\phi_{4}\partial_{\phi_{3}}+\phi_{3}\partial_{\phi_{4}}-\phi_{6}\partial_{\phi_{5}}+\phi_{5}\partial_{\phi_{6}}-\phi_{8}\partial_{\phi_{7}}+\phi_{7}\partial_{\phi_{8}}$ (4.44)
which again infers the symmetry algebra
$\mathfrak{u}(1)_{Y}=\operatorname{span}(X_{Y}),$ (4.45)
the Lie algebra of the hypercharge symmetry group $\mathsf{U}(1)_{Y}$. The only other possibility (4.43) corresponds to a 2-dimensional algebra, parametrized by the two now free parameters $b_{21},b_{78}$, since there are no restrictions on $b_{21},b_{78}$ in (4.43). Eqs. (4.28) and (2.6) then gives a symmetry algebra
$\mathfrak{u}(1)_{Y}\oplus\mathfrak{u}(1)_{\text{PQ}}=\operatorname{span}(X_{Y},X_{\text{PQ}}),$ (4.46)
where
$X_{\text{PQ}}=\phi_{2}\partial_{\phi_{1}}-\phi_{1}\partial_{\phi_{2}}+\phi_{4}\partial_{\phi_{3}}-\phi_{3}\partial_{\phi_{4}}-\phi_{6}\partial_{\phi_{5}}+\phi_{5}\partial_{\phi_{6}}-\phi_{8}\partial_{\phi_{7}}+\phi_{7}\partial_{\phi_{8}},$ (4.47)
which spans the Lie algebra $\mathfrak{u}(1)_{\text{PQ}}$ of the Peccei-Quinn $\mathsf{U}(1)$ symmetry. We may demonstrate this is a strict variational symmetry by applying the fact $\operatorname{pr}X_{\text{PQ}}(T)=0$ (i.e. the Peccei-Quinn symmetry is a SVS of the kinetic terms), together with Theorem 1. Alternatively show (2.21) holds for potential parameters (4.5) and (4.43), or just check the consistency with the results in [21] or e.g. Table 5 of [26].
4.3.2 The case $m_{11}^{2}=m_{22}^{2}$
We now assume
$m_{11}^{2}=m_{22}^{2},$ (4.48)
and substitute (4.48) in the determining equations (4.37). The 148 simplest of these equations are of the form (4.40) again, and we solve and substitute the resulting equations in the total system of determining equations. The only free parameters $b_{ij}$ in the determining equations can now be taken as
$B_{2}=\{b_{21},b_{25},b_{51},b_{78}\}.$ (4.49)
Because of (4.48), the mass-squared matrix will remain diagonal under any additional Higgs basis transformations, hence we have freedom to reduce the potential some more, without introducing a new $m_{12}^{2}$ parameter. We therefore choose to diagonalize the matrix $\Lambda$ of (4.17), which means
$\lambda_{6}=\lambda_{7},\quad\Im(\lambda_{5})=0,$ (4.50)
cf. (4.24), where the latter reduction $\Im(\lambda_{5})=0$ is the same as before, cf. (4.5).
Assuming $\Re(\lambda_{5})\neq 0$
Additionally, we will now assume
$\Re(\lambda_{5})\neq 0,$ (4.51)
since $\Re(\lambda_{5})=0$ would also have let us reduce the potential by eliminating $\Im(\lambda_{6})$. By substituting (4.50) into the determining equations and solving for the parameters (4.49), with the assumption (4.51), we get 3 different solutions for the parameters $B_{2}$, cf. (4.49), and hence 3 possible Lie symmetry algebras: Let
$H_{1}$ $=\phi_{6}\partial_{\phi_{1}}-\phi_{5}\partial_{\phi_{2}}+\phi_{8}\partial_{\phi_{3}}-\phi_{7}\partial_{\phi_{4}}+\phi_{2}\partial_{\phi_{5}}-\phi_{1}\partial_{\phi_{6}}+\phi_{4}\partial_{\phi_{7}}-\phi_{3}\partial_{\phi_{8}},$
$H_{2}$ $=\phi_{5}\partial_{\phi_{1}}+\phi_{6}\partial_{\phi_{2}}+\phi_{7}\partial_{\phi_{3}}+\phi_{8}\partial_{\phi_{4}}-\phi_{1}\partial_{\phi_{5}}-\phi_{2}\partial_{\phi_{6}}-\phi_{3}\partial_{\phi_{7}}-\phi_{4}\partial_{\phi_{8}}.$ (4.52)
Then the Lie algebra of the reparametrization group $\mathsf{SU}(2)$ (i.e. the basis transformations), denoted $\mathfrak{su}(2)_{\text{HF}}$ (where ”HF” stands for ”Higgs Family”), will be spanned by the latter generators and
$H_{3}\equiv X_{\text{PQ}},$ (4.53)
c.f. (4.47), namely
$\mathfrak{su}(2)_{\text{HF}}=\operatorname{span}(H_{1},H_{2},H_{3}).$ (4.54)
The Lie algebra in this basis has the commutation rules [^11] Remember we are employing the mathematicians’ definition of a Lie algebra, and are hence working in an anti-Hermitian basis.
$[H_{i},H_{j}]=-2\epsilon^{ijk}H_{k}.$ (4.55)
Now, the first solution for the $b$’s is
$\text{I}:\quad b_{25}$ $=0,\quad b_{78}=-b_{21},$
$\text{when}\quad\lambda_{2}$ $=\lambda_{3}+\lambda_{4}+\Re(\lambda_{5})\wedge\lambda_{1}=\lambda_{2}$ (4.56)
with $b_{51}$ free, and this holds for the displayed parameter conditions, of course in addition to the other conditions like (4.5). The two other solutions are
$\text{II}:\quad b_{51}=0,\quad b_{78}=-b_{21}$ $\text{III}:\quad b_{25}=b_{51}=0,\quad b_{78}=-b_{21}$ (4.57)
where both solutions hold for several, different conditions on the potential parameters, similar to the ones given in (4.3.2). Then the 3 symmetry algebras are respectively
$\mathfrak{g}_{\text{I}}$ $=\operatorname{span}(H_{2},X_{Y}),$ (4.58)
$\mathfrak{g}_{\text{II}}$ $=\operatorname{span}(H_{1},X_{Y}),$ (4.59)
$\mathfrak{u}(1)_{Y}$ $=\operatorname{span}(X_{Y}).$ (4.60)
We can conclude all these algebras are strict variational symmetries, without considering all the distinct sets of conditions on the parameters of the potential: The prolonged infinitesimal elements $\operatorname{pr}X$ of all generators $X$ in (4.58) will annihilate the kinetic terms when applied on the most general $\mathcal{L}$. This may be checked explicitly, but it also follows from the well-known fact that $\mathsf{SU}(2)_{\text{HF}}$ is a strict variational symmetry group of the kinetic terms of the 2HDM, see [34] for a proof that the strict variational symmetries of the kinetic terms are $\mathfrak{su}(N)$ for a general NHDM. Then, Theorem 1 guarantees these symmetries are strictly variational, independent of the conditions on the parameters of the potential. Moreover, we will in section 4.3.3 show that the symmetries given by $\mathfrak{g}_{\text{I}}$ and $\mathfrak{g}_{\text{II}}$ are equivalent, which means that there is a new Higgs basis where $\mathfrak{g}_{\text{I}}$ will be the same as $\mathfrak{g}_{\text{II}}$ was in the old basis.
Assuming $\Re(\lambda_{5})=0$
We now consider the special case where
$\Re(\lambda_{5})=0,$ (4.61)
in addition to (4.48) and (4.50). By a rephasing of the doublets, we may now eliminate $\Im(\lambda_{6})$,
$\Im(\lambda_{6})=0,$ (4.62)
without introducing new parameters.
Assuming $\Re(\lambda_{5})=0$ and $\lambda_{1}+\lambda_{2}\neq 2(\lambda_{3}+\lambda_{4})$
If $\lambda_{1}+\lambda_{2}=2(\lambda_{3}+\lambda_{4})$ the matrix $\Lambda$ will be proportional to identity, and we can perform $\mathsf{SO}(3)$ rotations without altering the elements of $\Lambda$. Then we may freely rotate $L$ while $M$ remains zero, such that $L$ points in a desired direction, e.g. in the $x$-direction, and then will $\lambda_{1}=\lambda_{2}$. Hence, we will first consider the case where
$\lambda_{1}+\lambda_{2}\neq 2(\lambda_{3}+\lambda_{4}),$ (4.63)
in addition to all the previous assumptions, like (4.61) and (4.62). By solving the determining equations for this case, we get 3 solutions for the surviving parameters $B_{2}$, cf. (4.49). In the same manner as in earlier sections, the resulting Lie symmetry algebras are
$\mathfrak{g}_{1}$ $=\operatorname{span}(H_{3},X_{Y}),$ (4.64)
$\mathfrak{u}(1)_{Y}$ $=\operatorname{span}(X_{Y}),$ (4.65)
where two different solutions for the $b$’s correspond to the same algebra $\mathfrak{u}(1)_{Y}$, while $\mathfrak{g}_{1}$ is isomorphic to to $\mathfrak{u}(1)\oplus\mathfrak{u}(1)_{Y}$. In the same manner as before, all symmetries given by (4.64) are, by Theorem 1, strictly variational, since they annihilate the kinetic terms of the 2HDM.
Assuming $\Re(\lambda_{5})=0$ and $\lambda_{1}+\lambda_{2}=2(\lambda_{3}+\lambda_{4})$
We now consider the last case with, as explained in the beginning of the last paragraph,
$\lambda_{1}+\lambda_{2}=2(\lambda_{3}+\lambda_{4})\wedge\lambda_{1}=\lambda_{2},$ (4.66)
in addition to (4.48), (4.50), (4.61) and (4.62). Under these conditions we get two solutions for the $b$’s. In the first solution, $\Re(\lambda_{6})$ is free, and the corresponding symmetry algebra is
$\mathfrak{g}_{1}=\operatorname{span}(H_{1},X_{Y}),$ (4.67)
which again is isomorphic to $\mathfrak{u}(1)\oplus\mathfrak{u}(1)_{Y}$. The second solution is that all the four parameters of $B_{2}$ are free, cf. (4.49), while $\Re(\lambda_{6})=0$. In this case the symmetry algebra is
$\mathfrak{su}(2)_{\text{HF}}\oplus\mathfrak{u}(1)_{Y}=\operatorname{span}(H_{1},H_{2},H_{3},X_{Y}).$ (4.68)
Both symmetry algebras correspond to strict variational symmetries, due to Theorem 1.
4.3.3 The inequivalent Lie point symmetries of the 2HDM
All the 2-dimensional Lie algebras found in the analysis are equivalent, since a rotation
$H_{1}\to H_{2}\to H_{3}\to H_{1}$ (4.69)
of $\mathfrak{su}(2)_{\text{HF}}$ generators will be an inner automorphism of $\mathfrak{su}(2)_{\text{HF}}$, which means it may be implemented by a matrix conjugation by a matrix $U\in\mathsf{SU}(2)$, i.e. $UH_{1}U^{\dagger}=H_{2}$, etc. Moreover, a conjugation by a $\mathsf{SU}(2)$ matrix $U$ corresponds to a change of Higgs doublet basis, which means all the 2-dimensional symmetries we found really are the same symmetry transformation, expressed in different Higgs bases. Here the other element $X_{Y}$ of the 2-dimensional algebras corresponds to a complex phase, and commutes with all other (series of) generators, including $U$, and is not affected by the automorphism. Finally, the rotation of generators (4.69) is an automorphism since it conserves the commutator, and it is inner because all automorphisms of $\mathfrak{su}(2)$ are inner (its Dynkin diagram is just a dot with no non-trivial diagram automorphisms). Hence we have shown there are only three possible, inequivalent scalar Lie point symmetry algebras of the 2HDM, namely
$\mathfrak{su}(2)\oplus\mathfrak{u}(1)_{Y},\quad\mathfrak{u}(1)\oplus\mathfrak{u}(1)_{Y},\quad\mathfrak{u}(1)_{Y}$ (4.70)
and they are all strictly variational due to Theorem 1, consistent with the results in [21, 32]. Gauge symmetries are not included in (4.70), since we only have considered (purely) scalar symmetries. We have in addition demonstrated something new, namely that there are no scalar divergence or scalar non-variational Lie point symmetries in the 2HDM, since Lie’s method finds all symmetries of systems of PDEs, cf. (2.17). The absence of divergence and non-variational Lie point symmetries in the 2HDM stands in contrast to the situation in singlet extensions of the SM, where such symmetries do appear [50].
Furthermore, we have tested the soundness of our implementation of Lie symmetry analysis for models with many parameters, variables and reparametrization freedom, by applying the method on a well-understood example and replicating well-known results. Of course, the derivation in [21] of the strict variational point symmetries of the 2HDM is more effective, transparent and elegant, and also includes discrete symmetries in the same proof. However, the method applied in this section has the advantage that it is universal, and that it detects all divergence and non-variational symmetries as well as strict variational symmetries. Lie symmetry analysis could, at least in principle, be applied to any model, including any NHDM, in contrast to methods only considering the possible symmetry transformations of gauge invariant, scalar bilinears, cf. (4.9). The method applied in this work also yields the exact parameter conditions for all symmetries, although this may not be an essential feature in a classification of symmetries.
5 Summary and outlook
We have demonstrated, by example, how Lie symmetry analysis of field equations can be applied to quantum physical models with many variables, parameters and reparametrization freedom. In our analysis of the two-Higgs-doublet model (2HDM), we reproduced the well-known strict variational Lie point symmetries of the model and also showed the model forbids divergence and non-variational Lie point symmetries. This observation is essential: Divergence symmetries are, under the usual boundary conditions, symmetries of the action and will, provided the measure of Feynman’s path integral is invariant, be preserved under quantization in the same manner as strict variational symmetries. Moreover, we proved three general results that may simplify Lie symmetry analysis for a wide class of particle physics models. The most relevant result for the present work was Theorem 1, which enables us to decide whether scalar Lie point symmetries are variational for a large class of theories, simply by considering the actions of the symmetries on the kinetic terms. This criterion is independent of the possibly numerous parameter conditions accompanying the different potentials that may realize the same symmetry. The two other results, Proposition 2.5 and Corollary 1, were not strictly necessary for the present analysis, but they may substantially simplify the computations of variational, scalar Lie point symmetries of other, larger models of the type NHDM+KS. This is done by reducing the spacetime dimension, and hence the number of spacetime variables and gauge fields, while getting exactly the same results for scalar variational symmetries. Lie symmetry analysis only detects continuous (Lie) symmetries. Nevertheless, any discrete symmetries not detected by Lie’s method, may be identified through considering the automorphism groups of the Lie symmetry algebras computed through Lie symmetry analysis. In this way, the maximal, faithfully acting variational symmetry group of a theory may be determined. We leave a systematic study of such automorphisms and their role in detecting discrete symmetries of large quantum physical models for future work.
Acknowledgments
The author is grateful to Stylianos Dimas for providing the latest version of SYM [45] and a Mathematica notebook demonstrating its use.
Appendix A A proof of Proposition 2.5
The most general NHDM+KS Lagrangian in $d$ spacetime dimensions was given by (2.50). We may expand the Lagrangian (2.50) in purely scalar terms and terms linear and quadratic in the gauge fields, as follows
$\mathcal{L}_{d}(x_{0},\ldots,x_{d-1})$ $=\partial_{\mu}\Phi_{n}^{\dagger}\partial^{\mu}\Phi_{n}+(\partial^{\mu}\Phi_{n}^{\dagger})G_{\mu}\Phi_{n}+\Phi_{n}^{\dagger}G_{\mu}^{\dagger}\partial^{\mu}\Phi_{n}+\Phi_{n}^{\dagger}G_{\mu}^{\dagger}G^{\mu}\Phi_{n}$
$+\frac{1}{2}\partial_{\mu}s_{m}\partial^{\mu}s_{m}-V(\Phi,s)+T_{\text{GB}},$ (A.1)
where
$G^{\mu}=ig(\sigma_{i}/2)W^{\mu}_{i}+ig^{\prime}YB^{\mu}$ (A.2)
represent the gauge bosons and where e.g. $\mu\in\{0,1,2,3\}$ for $\mathcal{L}_{4}$ while $\mu=0$ for $\mathcal{L}_{1}$. Moreover, $T_{\text{GB}}$ is the sum of the kinetic terms of the gauge bosons given by (2.51), which will be irrelevant for the proof of Proposition 2.5, since we only will be considering scalar symmetries. The relevant 1-prolongation of the scalar infinitesimal generator for the $d$-dimensional case is given by
$\operatorname{pr}^{(1)}X_{d}$ $=\sum_{\mu=0}^{d-1}\sum_{i,j=1}^{4N+K}(\eta_{i}\partial_{\phi_{i}}+\frac{\partial\phi_{j}}{\partial{x_{\mu}}}\frac{\partial\eta_{i}}{\partial\phi_{j}}\frac{\partial}{\partial(\partial\phi_{i}/\partial x_{\mu})}),$ (A.3)
where we identify
$\phi_{4N+k}\equiv s_{k},\quad\text{for}\;k\geq 1.$ (A.4)
[…]
Proof.
First we consider strict variational symmetries, i.e. symmetries with $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}=0$ (no sum over $d$), cf. (2.21) with $\xi=0$.
Denote $\partial_{\phi_{i}}\Phi_{j}\equiv\Phi_{j,i}$ and similar for Hermitian conjugated fields. Then e.g. $\Phi_{1,4}^{\dagger}=(0,-i)$. The effect of $\operatorname{pr}^{(1)}X_{d}$ of (A.3) on $\mathcal{L}_{d}$, given by (A), may be written
$\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}$ $=\eta_{i}(\partial^{\mu}\Phi_{n}^{\dagger})G_{\mu}\Phi_{n,i}+\eta_{i}\Phi_{n,i}^{\dagger}G_{\mu}^{\dagger}\partial^{\mu}\Phi_{n}+\eta_{i}\Phi_{n,i}^{\dagger}G_{\mu}^{\dagger}G^{\mu}\Phi_{n}+\eta_{i}\Phi_{n}^{\dagger}G_{\mu}^{\dagger}G^{\mu}\Phi_{n,i}$
$-\eta_{i}\frac{\partial V(\Phi,s)}{\partial\phi_{i}}+\frac{\partial\phi_{j}}{\partial{x_{\mu}}}\frac{\partial\eta_{i}}{\partial\phi_{j}}\frac{\partial\phi_{i}}{\partial x^{\mu}}+\frac{\partial\phi_{j}}{\partial{x_{\mu}}}\frac{\partial\eta_{i}}{\partial\phi_{j}}(\Phi_{n,i}^{\dagger}G_{\mu}\Phi_{n}+\Phi_{n}^{\dagger}G_{\mu}^{\dagger}\Phi_{n,i}+\delta_{m,(i-4N)}\partial_{\mu}s_{m})$
$=0,$ (A.5)
where $0\leq\mu\leq d-1$ and $1\leq K$. Assume that $\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1}=0$ holds, we would like to show that then $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}=0$ holds, with the same $\eta$’s. This will mean that a scalar symmetry of the $d=1$ theory also is a scalar symmetry of the theory with arbitrary dimension $d$. Now, $\eta_{i}{\partial_{\phi_{i}}V(\Phi)}{}=0$, since all other terms are proportional to gauge bosons $G^{\mu}$ or derivatives $\partial\phi_{j}/\partial x_{\mu}$ (which includes terms $\partial_{\mu}s_{m}$), and hence cannot cancel $\eta_{i}{\partial_{\phi_{i}}V(\Phi,s)}{}$. This infers that $\eta_{i}{\partial_{\phi_{i}}V(\Phi)}{}=0$ in $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}$ as well, since it is the same expression, when suppressing spacetime variables. The other terms all depend on $\mu$, and since they cancel for $\mu=0$ in $\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1}$, they also have to cancel for each $\mu$ in $\operatorname{pr}^{(1)}X_{4}\mathcal{L}_{4}$, since the structure of the terms is the same for all $\mu$.
Conversely, assume $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}=0$, that is, we have a symmetry of the theory of some fixed, but arbitrary dimension $d>1$. Then (A) holds where $\mu$ this time is a sum from 0 to d1d-1. We will show that then $\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1}=0$ for the same $\eta$’s, which means the $d=1$ theory has the same symmetry. Again, since all other terms are proportional to gauge bosons or derivatives, $\eta_{i}{\partial_{\phi_{i}}V(\Phi)}{}=0$, and the same is true in the $d=1$ theory. Moreover, the other terms have to cancel for each $\mu$ individually: This is so because terms involving gauge bosons $G^{\mu}$ cannot be canceled by terms involving gauge bosons $G^{\nu}$ for $\nu\neq\mu$, since they are independent fields. Moreover the term $\frac{\partial\phi_{j}}{\partial{x_{\mu}}}\frac{\partial\eta_{i}}{\partial\phi_{j}}\frac{\partial\phi_{i}}{\partial x^{\mu}}$ for one specific, fixed $\mu$ cannot be cancelled by a term of the same type with fixed spacetime index $\nu\neq\mu$, since $\eta_{i}$ doesn’t involve spacetime derivatives of fields. There are no other terms that can cancel these terms either. So terms containing $\mu$ in (A) cancel for each choice of $\mu$ separately, and hence they will cancel for $\mu=0$ in the $d=1$ theory as well, and the symmetry for the $d>1$ theory is also a symmetry for the $d=1$ theory. So the scalar, SVSs for $\mathcal{L}_{d}$ and $\mathcal{L}_{1}$ are the same, hence they are the same for $\mathcal{L}_{d}$ for all $d\in\mathbb{N}$.
We now turn to the case where $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}=\operatorname{Div}\beta\equiv d_{\mu}\beta^{\mu}$ for some local vector field $\beta^{\mu}(x,\phi,\phi_{,\alpha},\phi_{,\alpha\beta})$, that is not identically zero. The first order infinitesimal variation $\delta\mathcal{L}_{d}=\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}$ is a divergence if and only if the Euler operator $E$ applied on $\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d}$ is zero, cf. (2.26), which means
$E_{j}(\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d})=(\frac{\partial}{\partial y^{j}}-\frac{d}{dx^{\mu}}\frac{\partial}{\partial y^{j}_{,\mu}})(\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d})=0,$ (A.6)
for all $1\leq j\leq(4N+K+4d)$, where the dependent variables $y^{j}$ are the $4N+K$ Higgs field components and the $4d$ gauge field components, and where $d$ again is the number of independent variables (the dimension). First, assume $E(\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1})=0$, i.e. we have a divergence symmetry in the $d=1$ theory. The term $\eta_{i}{\partial_{\phi_{i}}V(\Phi,s)}{}$ only depends on undifferentiated scalars, hence if $E_{j}$ corresponds to a Higgs field $y^{j}=\phi_{j}$, then $E_{j}(\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1})=0$ infer $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))=0$. This is so because the term $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))$ cannot be canceled by any other term in $E_{j}(\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1})$ since they all will be proportional to either terms of the type $G_{\mu}$, $\partial_{\mu}\phi_{k}$ or a $\partial_{\mu}\partial^{\mu}\phi_{k}$ (all with $\mu=0$), in contrast to the term $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))$. Hence $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))=0$ also will be true in the $d>1$ theory, since these terms are the same for all $d$, suppressing spacetime variables. All the remaining terms of (A) have got two factors of the type $G_{\mu}$ or $\partial_{\mu}\phi_{k}$ for $\mu=0$, and no matter which $E_{i}$ we apply, at least one such factor survives in each non-annihilated term, so all these terms are tagged with the index $\mu=0$. Since the $d>1$ dimensional theory have $d$ copies of these terms, where each copy is tagged with an index $\mu$, $0\leq\mu\leq d-1$, each of the $d$ copies have to vanish separately, since they are structurally the same, only distinguished by the index $\mu$. Hence we also have $E(\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d})=0$.
Conversely, assume $E(\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d})=0$, i.e. we have a divergence symmetry in $d>1$ spacetime dimensions. By the same reason as before the terms $\partial_{\phi_{j}}(\eta_{i}\partial_{\phi_{i}}V(\Phi,s))=0$ in the $d>1$ theory, and hence the same is true in the $d=1$ theory, since those terms are the same for all choices of $d$. The remaining terms of $E(\operatorname{pr}^{(1)}X_{d}\mathcal{L}_{d})$ can be grouped in $d$ distinct, non-overlapping categories determined by the value of $\mu$ in factors of the types $G_{\mu}$, $\partial_{\mu}\phi_{k}$ or $\partial_{\check{\mu}}\partial^{\check{\mu}}\phi_{k}$. The terms in each category have to cancel separately, since all terms in one category are linearly independent of any other term in another category. [^12] Note that the the conditions for variational symmetries, including divergence symmetries, don’t presuppose that the Euler-Lagrange equations of the Lagrangian holds. Hence there is no relation between $\partial_{\check{\mu}}\partial^{\check{\mu}}\phi_{k}$ and $\partial_{\check{\nu}}\partial^{\check{\nu}}\phi_{k}$ for $\check{\mu}\neq\check{\nu}$. Hence the terms in the category $\mu=0$ cancel, and hence $E(\operatorname{pr}^{(1)}X_{1}\mathcal{L}_{1})=0$. Therefore, the scalar symmetries of $\mathcal{L}_{d}$ are the same for all $d$, since $\mathcal{L}_{d}$ have the same scalar symmetries as $\mathcal{L}_{1}$ for all $d$. Finally, we note that the arguments hold equally well if $K=0$, i.e. we have a pure NHDM. ∎
We don’t expect Proposition 2.5 to hold in general for non-variational symmetries, since the kinetic terms of the gauge bosons consist of cross terms of different spacetime derivatives, and these terms will now matter, since a symmetry of the Euler-Lagrange equations here infer that
$\operatorname{pr}^{(2)}X_{d}(E(\mathcal{L}_{d}))=0\quad\text{when}\quad E(\mathcal{L}_{d})=0,$ (A.7)
and the kinetic terms of the gauge bosons will be involved in the ”on-shell” condition $E(\mathcal{L}_{d})=0$, while this was not the case for the ”off-shell” (i.e. variational), scalar symmetries in Proposition 2.5. In contrast, the condition for a variational symmetry here is
$E(\operatorname{pr}^{(1)}X_{d}(\mathcal{L}_{d}))=0,$ (A.8)
with no assumption on the validity of the Euler-Lagrange equations. Moreover, for $d=1$ the kinetic terms of the gauge bosons are zero due to the antisymmetry of the spacetime indices, while they are non-zero for $d>1$, which induces an asymmetry between the $d=1$ and $d>1$ cases in the on-shell condition $E(\mathcal{L}_{d})=0$ for the two cases. Finally, the Euler-Lagrange equations will infer dependencies between derivatives like $\partial_{\check{\mu}}\partial^{\check{\mu}}\phi_{k}$ and $\partial_{\check{\nu}}\partial^{\check{\nu}}\phi_{k}$ for $\check{\mu}\neq\check{\nu}$, in contrast to the off-shell situation in Proposition 2.5, cf. footnote 12.
Appendix B Symmetries and changes of Higgs bases
To what extent does a change of Higgs basis change the symmetry of an equation or a Lagrangian $\mathcal{L}$, and how is it still detectable by Lie’s method? As an example, consider a linear transformation of the scalars given by a matrix $S$, which is a strict variational symmetry if
$\mathcal{L}(\phi)=\mathcal{L}(S\phi).$ (B.1)
If we then consider a Higgs basis shift ${\phi}^{\prime}=U\phi$, this possibly obfuscates a symmetry $S$, manifest in the unprimed basis $\phi$. The symmetry will still be present as an infinitesimal symmetry in the primed basis, although it may not be manifest (i.e. it is hidden). For if
$\mathcal{L}(\phi)=\mathcal{L}(U^{\dagger}\phi^{\prime})\equiv\mathcal{L}^{\prime}(\phi^{\prime}),$ (B.2)
then $S^{\prime}=USU^{\dagger}$ is a symmetry of the transformed Lagrangian $\mathcal{L}^{\prime}$
$\mathcal{L}^{\prime}(\phi^{\prime})=\mathcal{L}^{\prime}(S^{\prime}\phi^{\prime}),$ (B.3)
if $S$ is a symmetry of the original Lagrangian, since
$\mathcal{L}^{\prime}(S^{\prime}\phi^{\prime})=\mathcal{L}(U^{\dagger}USU^{\dagger}\phi^{\prime})=\mathcal{L}({\phi})=\mathcal{L}^{\prime}(\phi^{\prime}).$ (B.4)
If $S$ is an infinitesimal symmetry, $S^{\prime}$ is of course infinitesimal as well: For if $S=I+\epsilon X$ with $\epsilon$ an infinitesimal parameter, then $S^{\prime}=I+\epsilon UXU^{\dagger}$. Then the one-parameter group $\exp(tUXU^{\dagger})$, for $t\in\mathbb{R}$, is continuous and connected to the identity, and is hence a Lie symmetry that will be detected by Lie symmetry analysis, albeit not a manifest symmetry.